### v2

# CGRA4ML: A Framework to Implement Modern Neural Networks for Scientific Edge Computing

Abstract
Scientific edge computing increasingly relies on hardware-accelerated neural networks to implement complex, near-sensor processing at extremely high throughputs and low latencies. Existing frameworks like hls4ml are effective for smaller models, but struggle with larger, modern neural networks due to their requirement of spatially implementing the neural network layers and storing all weights in on-chip memory. CGRA4ML is an open-source, modular framework designed to bridge the gap between neural network model complexity and extreme performance requirements. CGRA4ML extends the capabilities of hls4ml by allowing off-chip data storage and supporting a broader range of neural network architectures, including models like ResNet, PointNet, and transformers. Unlike hls4ml, CGRA4ML generates SystemVerilog RTL, making it more suitable for targeting ASIC and FPGA design flows. We demonstrate the effectiveness of our framework by implementing and scaling larger models that were previously unattainable with hls4ml, showcasing its adaptability and efficiency in handling complex computations. CGRA4ML also introduces an extensive verification framework, with a generated runtime firmware that enables its integration into different SoC platforms. CGRA4ML’s minimal and modular infrastructure of Python API, SystemVerilog hardware, Tcl toolflows, and C runtime, facilitates easy integration and experimentation, allowing scientists to focus on innovation rather than the intricacies of hardware design and optimization.
I Introduction
Figure 1: CGRA4ML workflow as outlined in Sec. III. Users first build quantized neural networks and train them in a quantization-aware manner using qkeras (Sec. V-B). Users then define a CGRA definition, which generates vendor-agnostic SystemVerilog RTL hardware specifications and TCL tool flows (Sec. IV). The model can then be exported to generate weights and a C runtime firmware (Sec. V-C). The generated hardware IP is then verified comprehensively with the model and firmware (Sec. V-D) using our randomized, transactional SystemVerilog testbench suite with DPI-C extensions. Finally, the bitstream generated from the FPGA toolflow and the C firmware can be loaded into an FPGA to be tested in seconds (Sec. V-E1). After such rapid prototyping, the same hardware design can be moved to ASIC (Sec. V-E2).
Scientific applications have unique and challenging constraints to process the data quickly and closely to where the data is generated [1]. This enforces strict real-time requirements to process the information at tremendously high throughputs (TB/s) and low latencies. Thus, fast machine-learning scientific applications often require hardware accelerators implemented on FPGAs or ASICs to meet the stringent performance requirements [2].
hls4ml is a popular tool for developing hardware-accelerated scientific neural networks, which implements a dataflow-style architecture with layer-specific datapaths and on-chip weight storage [3]. While this methodology provides high throughput and low latency, it does not support modern, deep neural network models (DNNs) that are too big to fit fully on-chip, requiring off-chip movement of weights and data. In addition, modern neural network models require the development of optimized layer architectures, e.g., a transformer cannot easily be implemented in hls4ml due to its intensive customization and large memory requirements for weights. Most accelerators presented in the literature that support DNNs do not offer usable frameworks with parametrization of their architecture or provide support for verification and SoC integration, as described in Sec. V. The frameworks available for accelerating larger models in FPGAs do not fulfill the requirements of the scientific computing community, such as more robust quantization, extensibility to add more features, and the need to eventually move from FPGA to ASIC implementation, as summarized in Table I.
CGRA4ML bridges the gap between hls4ml and the needs of the scientific community to implement modern, large neural network models using an easy-to-use, high-level workflow. Like hls4ml, CGRA4ML is entirely open source, allowing the scientific community to use, develop, and build upon it. CGRA4ML is easily programmable and scalable, allowing seamless scaling to fit the resources at hand while being easy and intuitive to integrate into the fast design prototyping flow. CGRA4ML generates a parameterizable Coarse-Grained Reconfigurable Array (CGRA) to facilitate the spatial reuse of processing elements and leverage off-chip data storage to deploy modern neural network features. Our Python API works on top of qkeras, a popular library for quantizing and training neural networks. CGRA4ML emits RTL instead of HLS code, making it more amenable to ASIC and FPGA design flows compared to hls4ml.
CGRA4ML builds and configures parametrized flexible CGRA that is then programmed to compute different neural network models. CGRA4ML facilitates resource sharing across neural network layers, as opposed to hls4ml, which implements each layer as a separate datapath. CGRA4ML scales beyond hls4ml capabilities, e.g., implementing larger ResNet models that fail to synthesize with hls4ml. This enables scientific applications that require more complex models that cannot fit within an FPGA or given area of an ASIC. We demonstrate how CGRA4ML can be extended to handle different neural network architectures, including ResNets, PointNet, and transformers. CGRA4ML implements its hardware as fully parametrized SystemVerilog RTL modules; the same implementation is supported in different FPGA and ASIC tools.
CGRA4ML is targeted to two user groups: (1) Scientists, embedded / IoT researchers who want to build and implement modern neural networks that are infeasible in hls4ml due to their size and complexity. (2) Hardware system designers who want to iteratively test new optimizations such as new datatypes, dataflow, and compression techniques.
CGRA4ML provides a Python API for the end-to-end workflow to allow scientists to build and implement their designs to production-ready quality without spending months on designing and verifying hardware, optimizing computation and data bottlenecks, integrating into an SoC, and writing, verifying, and optimizing runtime firmware (see Section III). The CGRA4ML infrastructure, shown in Fig 1 and described in Section V, is modular and extensible to allow the hardware researchers to modify any part, implement their optimizations, and get it working without building the entire system and optimizing every part of it themselves.
To demonstrate the capabilities of CGRA4ML and to compare with other tools our scientific users are familiar with, we present the implementations of ResNet-50, PointNet, and Particle Transformer models in Section VI. We also present the performance, power, and area comparisons of the FPGA and ASIC implementations generated by CGRA4ML.
The major contributions of CGRA4ML are:
  • A hardware + software + firmware co-design framework to specify, train, quantize, and implement diverse neural networks on FPGA and ASIC as CGRA-powered SoCs.
  • A CGRA architecture with dynamic reconfiguration, unified dataflow, and an efficient PE design to maximize hardware utilization for modern neural networks.
  • RTL-based high-performance, modular SoC design with open-source DMAs and AXI-compliant interfaces to support vendor-agnostic FPGA and ASIC design flows.
  • Automated firmware generation enabling hardware/ software partitioning that easily integrates the CGRA with on-chip processors. Our results show how to partition DNN layers across an ARM core and the CGRA.
  • Validation of results on AMD Xilinx FPGAs and ASIC implementation using Cadence tools to be taped out as an SoC with ARM-based NanoSoC platform.
  • Extensive verification support to test the neural network model against the generated RTL design and C runtime through randomized transactional testbenches.
II Background and Related Work
II-A Neural Networks for Scientific Computing
Machine learning is fueling scientific discoveries in particle physics, materials, dark matter, cosmology, nuclear physics, biomedical engineering, and health monitoring [2]. Specific examples include bio-signal classification [4], tracking magnetohydrodynamic (MHD) instabilities in fusion reactors [5], reinforcement learning for accelerator beam control [6], pulse detection for anti-neutrino detection [7], and accelerator controls for beam loss deblending [8].
Scientific computing increasingly relies on neural networks to process data with requirements for extremely low latencies and high throughputs. For example, the Large Hadron Collider (LHC) has $\mathcal{O}(100M)$ individual sensors generating data for each of the 40 million proton beam collisions per second [9]. Neural network-based autoencoders implemented as ASICs compress data to move it off chip [10]. LHC trigger systems filter events in real-time with sub-microsecond latency requirements and 40 MHz throughput requirements using FPGAs [11]
The scientific community has benefited greatly from the ease of hardware acceleration. Among the multiple tools [12] that allow FPGA implementation of neural networks, hls4ml is very popular [13]. Yet, hls4ml has drawbacks as discussed in the next section. CGRA4ML allows more complex models to be implemented on FPGA and ASIC-based SoCs, moving more computation to the edge. This enables more sophisticated neural networks to be deployed in scientific equipment.
II-B hls4ml
hls4ml is a Python library developed by physicists for machine learning inference in FPGAs [13]. It primarily supports models built and trained using QKeras, a library from Google for low-precision machine learning [14]. hls4ml converts the model, layer by layer into customized high-level synthesis (HLS) hardware blocks, as shown in Figure 2. hls4ml uses different HLS implementations for different HLS tools to generate the hardware accelerator depending on the resulting target (AMD Xilinx FPGA, Altera FPGA, ASIC, etc.). The most common target is AMD Xilinx Vitis/Vivado HLS, though other backends exist in various forms (Intel HLS Compiler, Siemens Catapult HLS, etc.). The dependence on HLS restricts the user to the supported vendors and requires HLS tool-specific reimplementation of each hardware layer, complicating the maintenance and verification. hls4ml’s support for ASIC implementation is primarily through Catapult HLS, which currently only supports a limited subset of layers.
The most common hls4ml design flow is to generate an overall IP for the entire neural network and then integrate that IP into the SoC using vendor-specific IPs such as DMAs. There are efforts for better end-to-end development. For example, hls4ml provides a PYNQ overlay, a Python-based runtime for AMD Xilinx Zynq FPGAs. This workflow allows the physicists to focus on building and training low-precision models for their custom applications without spending months on hardware design, implementation, and verification. Yet, by and large, the ML IP core integration for production-ready designs is a separate process from hls4ml.
hls4ml implements the neural network layer by layer, as shown in Fig. 2. The user controls reuse factor per layer, which defines the number of times a multiplier/accumulator is reused. The resources needed to synthesize a layer and the number of clock cycles before its hardware block can accept new data both depend on the reuse factor. Commonly, the same reuse factor is used for all layers. This results in heavy underutilization since all hardware blocks must wait for the slowest to complete. Manual tuning of the reuse factors per layer on each target FPGA to balance the dataflow takes a lot of time and effort.
While the layer-by-layer approach can offer in low-latency, efficient implementation of small models that fully fit into FPGAs, it does not support modern neural network models like transformers. There have been herculean efforts to implement models like ResNet-50 on up to 12 interconnected FPGAs [15]. Other large models, e.g., jet tagging using a transformer [9], are implemented on GPUs because they are impossible to implement with hls4ml. The data filtering pipeline from the detectors of the Large Hadron Collider at CERN implements a tiny autoencoder NN on an ASIC closest to the detector. The data from the ASIC is then processed by progressively more complex NNs on FPGAs and servers. CGRA4ML strives to provide a user-friendly, more customizable workflow similar to hls4ml while supporting larger DNNs, allowing more computations to be moved to the edge and improving the accuracy of the data filtering pipeline.
Figure 2: Layer-by-layer spatial implementation of neural networks using hls4ml. hls4ml offers multiple backends, one per family of devices. Each backend is a collection of layers defined as templated HLS. Each neural network layer is implemented as a separate datapath. Xilinx DMA IPs are connected to the HLS4ML design to move data in and out.
II-C CGRA
CGRAs were originally envisioned as coarse-grained FPGAs whose programmable elements work at the word-level instead of FPGA bit-level programming [16]. Early CGRAs were classified based on their integration into the processor. Tightly coupled CGRAs integrate into a processor data path and are executed as a custom instruction, e.g., Chess [17], MATRIX [18], and DySer [19]. Loosely coupled CGRAs act more as an accelerator that executes alongside the processor, executing in tandem with the processor and communicating via on-chip interconnect. Examples of loosely-coupled CGRAs include PipeRench [20], MorphoSys [21], CHARM [22], and FPCA [23]. CGRA4ML can be characterized as a loosely coupled CGRA as it implements sub-tasks of the neural network alongside a CPU with a focus on enabling hardware/software partitioning across the processor and processor.
CGRAs have seen a recent resurgence in industry and academic projects. Commercial CGRAs include Samsung Reconfigurable Processor [24] and the Renesas STP Reconfigurable Processor [25]. CGRA-ME [26] is an exemplary academic framework that supports a subset of CGRA designs.
CGRA4ML is targeted to optimize the neural networks for scientific applications with high throughput, low latency, and real-time requirements. As a result, the default CGRA design of CGRA4ML features small processing elements (PEs) to minimize resource use and an on-the-fly reconfiguration mechanism to reduce the reconfiguration overhead.
This hls4ml VTA FINN Vitis AI DLP OpenVino Caffeine CGRA-ME FlexCNN
work [13] [27] [28] [29] [30] [31] [32] [26] [33]
Open-source
Sub-8-bit quantization support
Larger neural networks support
Xilinx FPGA support
Intel FPGA support $\sim$
ASIC Implementation $\sim$
Basic verification
Generate production-ready runtime
Verification with runtime
TABLE I: Comparing CGRA4ML with existing ML to FPGA/ASIC frameworks.
II-D DNN Reconfigurable Architectures
Hundreds of accelerator architectures have been designed to accelerate deep neural networks in the past decade [34, 35, 36]. These accelerators can be categorized by their data reuse pattern as weight, input, output & row stationary and by their runtime flexibility into systolic arrays, CGRAs, and microcode processors.
We compare the CGRA4ML with state-of-the-art architectures. Caffeine [32] offers an HLS systolic array to process CNNs from the Caffe ML framework using weight major and input major mappings. Eyeriss [37] is an energy-efficient ASIC designed to accelerate CNNs using a 2D array of 168 fairly complex processing elements, each with a 16-bit multiply-accumulate, scratchpad, and a controller. ShiDianNao [38] uses a 2D mesh of functional units optimized towards the 2D feature maps of convolutional layers. Each processing element executes multiplications, additions, and comparisons using 16-bit fixed point arithmetic. The kernel elements are shifted right to left and up to down and accumulate locally.
While systolic arrays with simple PEs are optimal for matrix multiplications, they are underutilized and incur a significant overhead when computing more complex layers. CGRAs that allow greater flexibility in data routing optimally utilize their PEs but are able to fit fewer PEs within given resource constraints due to the complexity of the PEs, local scratchpad memories, and routing logic. CGRA4ML aims to strike a balance between these by providing a default CGRA engine with an efficient PE design to maximize PE utilization while allowing flexibility to dynamically regroup to process layers such as convolution maximally exploiting data reuse on multiple levels (see Section IV). CGRA4ML is primarily output stationary; weights and input pixels are stored in on-chip memories to maximize utilization.
II-E ML to FPGA/ASIC Frameworks
There are a few end-to-end frameworks available for hardware implementation of neural networks [12]. DNNBuilder [39] and finn [28], implement a given model as a pipeline of layers, similar to hls4ml. DNNBuilder allows the users to customize two kinds of reuse factors: channel and kernel. finn takes in an ONNX model, possibly exported from Brevitas [40], to generate Vivado HLS IP. This can be verified in simulation before implementing on AMD Xilinx FPGAs. FINN provides a PYNQ driver for prototyping.
AMD Xilinx Vitis AI [29] is a closed-source library that implements a Deep Learning Processing Unit - an 8-bit micro-coded processor to process neural networks optimized through their stack on AMD Xilinx FPGAs. OpenVino is a similar stack for Intel FPGAs. Apache TVM, an open-source framework for embedded AI, implements a Versatile Tensor Accelerator (VTA) [27] as a GEMM processor using Xilinx HLS. LeFlow [41] emits HLS code, which has similar advantages and disadvantages as hls4ml.
While each of these frameworks targets different groups of users, the limitations in their approaches make them incompatible with the requirements of the scientific computing community, as listed in Table I. The popularity of hls4ml and its features: arbitrary quantization, Xilinx and Intel FPGA backends, and support for limited verification, demonstrate the unique needs of the scientific edge community. The models used in scientific applications require quantization to arbitrary bit-widths, which is made possible by qkeras, and not supported by Vitis AI, OpenVino, or CGRA-ME, among others. In addition, their workflow involves implementing their models on FPGAs and later moving to ASIC designs, which is not possible with Vitis AI, Apache VTA, and OpenVino since their backends are implemented in vendor-specific HLS. CGRA4ML fills this gap by providing a high-level API where the users can first configure a CGRA and export it as SystemVerilog RTL for ASIC or FPGA implementation from any vendor, then build models quantized to arbitrary bitwidths, train them using qkeras, and export the model runtime to be executed in any host CPU.
III CGRA4ML Overview
CGRA4ML is designed to satisfy the unique requirements of the scientific edge computing community, borrowing from the distinctive strengths of hls4ml and improving upon its limitations. This section describes the CGRA4ML workflow, as shown in Fig. 1 for implementing scientific computing applications, while Sec. V explains the technical details.
A typical scientific computing workflow involves building custom models with quantization-aware training. CGRA4ML’s Python API supports this workflow in the same way as hls4ml. qkeras performs quantization-aware training. qkeras is a library that offers robust quantization options and training algorithms [14]. After training, the user exports the model, which generates intermediate inputs and outputs as integer versions of the fixed-point representation for testing.
After defining the QKeras model, it is lowered into a list of bundles. A bundle is subclass of a qkeras layer that exists in the CGRA4ML library. Its C runtime defines its functionality and runtime parameters that map the QKeras layer to the CGRA. Bundles are further described in Sec. V-A.
CGRA4ML implements large and diverse neural networks by creating an efficient yet programmable computational array that can easily be shared across multiple neural network layers. The array’s computation and data movement are design-time parameterizable (Sec. IV-A) and run-time programmable (Sec. IV-B) to accommodate a variety of DNN layer types. CGRA4ML focuses on flexibility and data movement as a key enabler for efficient scientific neural network hardware acceleration. The CGRA and the system around it is implemented as SystemVerilog RTL, to target any FPGA or ASIC flow. This fundamentally differs from hls4ml, which emits HLS code that requires a separate HLS tool to lower its output to an RTL description.
CGRA4ML provides toolflows for ASIC implementation using Cadence and Synopsys tools. Users can add paths for their technology PDKs and memory compilers to the scripts and invoke them to synthesize place and route, check DRC violations, verify layout vs. schematic (LVS), and export GDSII files. CGRA4ML allows netlist verification using gate-level simulation. Users get power, performance, and area estimates through the generated reports. ASIC designers can edit and fine-tune the base scripts generated by CGRA4ML before the final tape-out. A system on chip generated through this workflow is being taped out with the ARM-based open-source NanoSoC platform as a proof of concept, as described in Sec. V-E2.
After the CGRA4ML design is implemented, it must be integrated into an SoC. CGRA4ML generates an AXI-compliant [42] high-performance IP powered by built-in open-source, design tool-agnostic AXI DMAs that can be integrated into different SoCs. CGRA4ML also directly generates the C firmware. The firmware is designed as a simple header-only library, allowing users to get the design working within minutes and then extend it by building the system-on-chip around it, adding other hardware for co-processing or moving data through external memory, ethernet, PCIe, etc. The firmware configures the CGRA and receives data to post-process and store in buffers for processing by the next layer. The workload is shared between the CGRA and the host CPU as described in Section V-C.
Verification dominates design time [43]. Thus, ensuring that the CGRA4ML implementation matches the neural network model is critical. CGRA4ML verification infrastructure can be invoked from Python API to target Verilator, AMD Xilinx Xsim, or Synopsys VCS simulators. This verifies the user’s neural network model, the SystemVerilog RTL of the generated CGRA hardware, and the generated C runtime together through a randomized SystemVerilog testbench and reports the error between the outputs of the user-defined Python model and the generated hardware.
IV CGRA4ML Hardware Architecture
CGRA4ML is a parametrizable, efficient CGRA optimized to execute complex neural network models. CGRA4ML uses on-chip memory to cache and reuse weights, reducing data movement and energy, a pixel shifter to facilitate data reuse for intermediate results, three high-performance, open-source AXI DMAs that move data in and out of the engine, and a hardware-based DMA controller with an AXI-Lite configuration. The engine is described as a vendor-agnostic SystemVerilog design. CGRA4ML generates the C firmware to control the engine and perform data post-processing. Figure 3 gives an overview of the CGRA architecture, interfaces, and firmware.
Figure 3: CGRA4ML architecture generated for AMD Xilinx Zynq SoC: Our CGRA-based engine (a) features one AXI-Lite port for configuration and three 128-bit AXI Manager ports for data movement. The CGRA engine (b) performs compute-heavy tasks. The engine has parameterizable bit widths, weights cache, and processing elements (PE) to fit the requirements of different models and meet hardware resource and performance constraints. CGRA4ML emits software for the ARM processor (c) to perform pixel-wise operations and handle the edge cases. .
The CGRA4ML engine features three high-bandwidth AXI-4 manager ports (up to 128-bit wide) to move inputs, weights, and outputs, and one AXI-Lite subordinate port to configure the engine with runtime parameters (see Table  II). Three AXI DMAs operate asynchronously to pull input and weight data from off-chip memory, feed them as AXI streams, receive the CGRA outputs as an AXI stream, and store them in on-chip memory. The control generator provides the configuration bits to the weights cache and pixel shifter, which process the configuration into a few bits and append them to the AXI-Stream data moving through the CGRA. CGRA4ML uses the industry-standard AMBA AXI interfaces, which enhance portability and ease the integration of the CGRA into the SoC.
IV-A Parameterized CGRA Engine Definition and Generation
A user sets the CGRA engine parameters using the Python API, which includes attributes of CGRA specification:
  • Number of Rows & Columns of PEs in the CGRA
  • Depth of weights cache
  • Bitwidth of inputs, kernel, outputs, and bias
  • Bitwidth of AXI interfaces (up to 128-bits)
  • Frequency, I/O delays
  • Max. batch, kernel sizes, in-channels, height, and width
  • Valid and Ready probabilities for randomized verification
The specification defines the CGRA architecture, which can be reprogrammed at runtime within these constraints to execute many different models. These attributes are parameterized into the SystemVerilog hardware to generate different CGRA engines quickly. The parameters are made static before synthesis based on the user-defined attributes.
The rows and columns dictate the number of PEs in the CGRA, which defines the hardware resource usage and the peak performance. The depth of the weights cache controls the weight reuse factor. These attributes provide a trade-off between performance and available hardware resources. While any model can be run on any specification, the hardware parameters can be optimized for specific models as described in Subsection IV-C. The bitwidths are customizable across the inputs, outputs, kernel, and bias. The PE datapaths are automatically adapted to match the requirements of these data types. Quantization is critical to meet high throughput and low latency requirements in scientific applications; thus, CGRA4ML supports fixed-point datatypes. The AXI interfaces are parameterized to easily match the SoC interfaces. The maximum limits of layer dimensions determine the register sizes within the hardware (see Table II).
CGRA4ML provides a robust verification environment using the randomized transactional testbench suite. The valid and ready probabilities define the randomness in toggling the s_valid and m_ready signals of the AXI interfaces. Setting the probabilities to 1 gives the fastest simulation, where data is transferred every cycle. This is helpful as a smoke test when a user intends to modify the hardware and debug their modification. Setting it to lower values, e.g., 0.1 and 0.01, stress tests the hardware using randomized transactions, exposing hardware bugs that smoke tests cannot reveal.
Fig. 3 (d) shows the CGRA4ML PEs, which include a multiplier, accumulator, two registers, and two multiplexers. Simplicity is favored over complexity to increase computational density and efficiency. CGRAs designed with a higher degree of flexibility in data movement and operation types suffer from the complexity required to program them and often utilize their resources less effectively. Therefore, our lightweight CGRA design aims to offer minimal programming complexity while supporting high utilization required for scientific edge computing. The PE columns can group themselves and process convolution kernels, sharing data to maximize data reuse. The pixel shifter maximizes data reuse across vertical convolution while maintaining modularity. The weights cache consists of two on-chip RAMs with parameterized depths in a ping-pong configuration to allow full throughput.
TABLE II: Runtime parameters of the unified dataflow
Description Array Partition
Conv Matmul Slice (S) Iterations (T)
$K_{H}$ Kernel Height 1
$K_{W}$ Kernel Width 1
$N$ Conv Batch 1
$W$ In Im. Width 1
$H$ In Im. Height I/O Rows $H_{S}{=}R$ $H_{T}{=}\lceil H/H_{S}\rceil$
$I$ In Channels In Cols $I_{S}{=}\lceil D_{W}/K_{H}\rceil$ $I_{T}{=}\lceil I/I_{S}\rceil$
$O$ Out Channels Out Cols $O_{S}{=}\lfloor C/K_{W}\rfloor$ $O_{T}{=}\lceil O/O_{S}\rceil$
  • $D_{W}{=}$ Depth of SRAM in weights rotator
  • $R,C{=}$ Rows & columns of PE array
Figure 4: Tiling: An output matrix with a shape $[H,O]$ is generated as underlineSlices of shape $[H_{S},O_{S}]$ (a), in $HT{\times}OT$ iTerations by a CGRA. The compile-time hardware parameters: depth of weights cache, rows & columns of CGRA ($R{\times}C$) affect the runtime parameters as shown in (b).
$X$ $:[N,H,W,I]$ input
$X_{S}$ $:[N,(H_{T},H_{S}),W,(I_{T},I_{S})]$ slicing
$X_{T}$ $:\underbrace{[I_{T},N,H_{T},W,I_{S},}_{\text{data beats}}\underbrace{H_{S}{+}K_{H}{-}1]}_{\text{parallel words}}$ reordering
$K$ $:[K_{H},K_{W},I,O]$ weights
$K_{S}$ $:[K_{H},K_{W},(I_{T},I_{S}),(O_{T},O_{S})]$ slicing
$K_{T}$ $:\underbrace{[I_{T},O_{T},}_{\text{(AXI-S Packets)}}\underbrace{I_{S},K_{H},}_{\text{(W-SRAM rows)}}\underbrace{O_{S},K_{W}]}_{\text{(parallel words)}}$ reordering
$Y$ $:[N,H,W,O]$ output
$Y_{S}$ $:[N,(H_{T},H_{S}),W,(O_{T},O_{S})]$ slicing
$Y_{T}$ $:\underbrace{[I_{T},O_{T},N,H_{T},W,}_{\text{data beats}}\underbrace{O_{S},H_{S}]}_{\text{parallel words}}$ reordering
$A^{\prime}$ $:[C,R]$ Accumulators of PE array
$A$ $:[O_{S},K_{W},R]$ dynamic regrouping
$A[:,:,:]\leftarrow 0$ for $i_{t}<I_{T}(=I/I_{S})$ do
     for $o_{t}<O_{T}(=O/O_{S})$ do
         for $n<N$ do
             for $h_{t}<H_{T}(=H/H_{S})$ do
                 for $w<W$ do
                      for $i_{s}<I_{S}$ do
                          for $k_{h}<K_{H}$ do
                              parallel for $o_{s}<O_{S}(=C/K_{W})$ do
                                  parallel for $k_{w}<K_{W}$ do
                                      parallel for $h_{s}<H_{S}(=R)$ do
                                          $x\leftarrow X_{T}[i_{t},n,h_{t},w,i_{n},h_{s}{+}k_{h}]$
                                          $k\leftarrow K_{T}[i_{t},o_{t},i_{s},k_{h},o_{s},k_{w}]$
                                          $A[o_{s},k_{w},h_{s}]\mathrel{{+}{=}}k*x$
1
2
3
4
5
                                          $Y[i_{t},o_{t},b,h_{t},w,:,:]\leftarrow A[:,{-}1,:]$
                                          parallel for $k_{w}<K_{W}$ do
                                              $A[:,k_{w},:]\leftarrow(k_{w}\equiv 0)$ $0:A[:,k_{w}{-}1,:]$
6
7
8
9
Algorithm 1 Unified Dataflow
IV-B Dynamic Reconfiguration and Dataflow
CGRA4ML’s dataflow and tiling pattern ensures weights are maximally reused, vastly reducing off-chip data movement. The dataflow aims to optimize data reuse patterns; the weights, inputs, and output data move through the CGRA in a predictable pattern for convolutions, dense layers, and matrix multiplications, and can be modified using the CGRA4ML attributes. This dataflow is automated, and its control is decentralized across the PEs.
The Python API generates the configuration bits for a layer and passes them to the C firmware. During the setup phase, the firmware writes attributes to the AXI-Lite configuration registers. During the operation, the DMA controller reads the configuration bytes and passes them along with data through the AXI DMAs. The Weights Cache and Pixel Shifter receive these configuration bits, process them into fewer bits, and pass them along with the AXI-Streams to the CGRA. The multiplexers inside the PEs respond to these control bits and reroute the data dynamically at every clock cycle. Since the configuration moves with the data, the CGRA PEs reconfigure themselves on the fly, grouping and sharing partial data to process a layer efficiently. The runtime parameters and the precise description of our default dataflow through the CGRA are presented in Table II and Algorithm 1, respectively.
Next, we describe the two main dataflow cases: Matrix Multiplication and Convolution. The flexibility and modularity of the CGRA4ML framework enables additional use cases with minimal modification. Additional layers and activation functions are added by modifying the C firmware.
IV-B1 Matrix Multiplication
Figure II shows the tiling (array partitioning), which is primarily output-stationary dataflow. When multiplying two matrices of sizes $[H,I]$ and $[I,O]$ to obtain an output matrix of size $[H,O]$, the output matrix is divided into slices of size $[H_{S},O_{S}]$ to be processed in $H_{T}{\times}O_{T}$ iterations. Each slice is processed in parallel using $R{\times}C{=}H_{S}{\times}O_{S}$ processing elements. When processing dense layers, the image batch is presented along $H$ to reuse weights across the batch.
IV-B2 Convolution
The unified dataflow is designed to maximize convolution data reuse. $R$ rows of processing elements compute $R{=}H_{S}$ Height slices. The pixel shifter moves $R{+}K_{H}{-}1$ values from the AXI DMA $K_{H}$ times to exploit data locality in vertical convolution. $C$ columns of PEs re-group dynamically into $O_{S}{=}\lfloor C/K_{W}\rfloor$ groups to process one Out-Channel slice $O_{S}$, while each such group of $K_{W}$ columns process the horizontal convolution, exploiting its data locality. $H_{S}{\times}O_{S}$ output pixels are computed in parallel by the $R{\times}C$ array of processing elements. The weights cache is $C$ words wide and $I_{S}K_{H}$ rows deep. The weights stored in the weights cache are rotated and reused $O_{T}NH_{T}W$ times, maximizing the weights reuse.
IV-C Performance Analysis
The number of clock cycles and the off-chip data movement required to process a layer on a CGRA with $R,C$ rows & columns of PEs are as follows:
Clock cycles $=O_{T}I_{T}(1+NH_{T}W(1+I_{S}K_{H}))$ (1)
Weight words $=O_{T}I_{T}I_{S}K_{H}C$ (2)
Input words $=O_{T}I_{T}NH_{T}WI_{S}(R+K_{H}/2)$ (3)
Output words $=NH_{\textbf{out}}W_{\textbf{out}}O.$ (4)
Increasing the number of PEs improves the performance by increasing the on-chip computation and reducing the number of iterations $O_{T},H_{T}$. The ratio between peak performance ($RC{\times}\text{Frequency}$) and real performance (MAC operations in a layer/time) depends on the ratio of PEs that are idle when computing a layer as follows:
Idle PE cols ratio ${=}[C\%K_{W}]/C{+}[O\%O_{S}]K_{W}/[CO_{T}]$ (5)
Idle PE rows ratio ${=}[H{\%}R]/H$ (6)
Expression (5) is minimized by our hardware and unified dataflow design that enables the $C$ columns of CGRA to dynamically group and process Out-Channels $O$. For example, for a CGRA with $C{=}96$, the ratio of idle columns in each iteration: $(C\%K_{W})/C$ is zero for the most common layers with $K_{W}{=}\{1,3\}$, and $1\%,5\%,8\%$ for less common layers with $K_{W}{=}\{5,7,11\}$. To further minimize the overall idle ratio of columns (Exp. 5), the user can pick $C$ such that $\lfloor C/K_{W}\rfloor$ is a factor of $O$ for most layers. For ResNets, multiples of 3 and 4 ($C{=}12,24,96$), and for Transformers, multiples of 8 ($C{=}8,16,32$) are such optimal choices. Among the $R$ columns of the CGRA, Expression 6 is minimized by choosing $R$ as the common factor of the $H$ of most layers of the DNNs the user wishes to run on the generated hardware. Figure 8 demonstrates the performance efficiency and off-chip data movement when processing ResNet50 on PE arrays of different sizes.
V CGRA4ML Framework
Figure 5: Modular and extensible infrastructure of the CGRA4ML framework as described in Section V. The Python API built around Google’s qkeras, extracts the fixed-point representations of intermediate tensors, generating the binary weights and model specification. The Python frontend also generates vendor-agnostic, synthesizable SystemVerilog RTL design of the CGRA per user specifications, optimized C runtime firmware and TCL toolflows for ASIC and FPGA implementations. Our comprehensive verification suite powered by randomized, transactional, SystemVerilog testbenches interfacing with the C firmware via DPI-C, then verifies the user’s model, the generated CGRA, and runtime together as it would run on an embedded system, emulating cases like memory congestion. The CI/CD pipeline tests the entire framework on a set of models over different CGRA configurations
CGRA4ML is a modular, extensible framework with a Python frontend that processes the neural network models and generates SystemVerilog RTL of specified CGRA hardware and SoC components, TCL toolflows for FPGA and ASIC design, a production-ready C runtime firmware, and the SystemVerilog testbench suite that verifies the users’ model, generated hardware and C runtime comprehensively. This section describes the technical details of our infrastructure outlined in Fig. 5.
V-A Bundles
The Bundle is the core building block of the CGRA4ML infrastructure that acts as the gateway between the neural network built by the user and the version that runs on the hardware. In the Python front-end, Bundle is implemented as a subclass of qkeras Layer with additional functions. In the C runtime, Bundle is implemented as a set of runtime parameters and C functions that operate on them. A bundle is a set of qkeras layers that can be deterministically executed on our hardware framework, as seen in Figure 6. The layers inside a bundle can be enabled or disabled, parametrized by a set of runtime parameters, and skip connections can be optionally added. A list of such Bundles forms various architectures the scientific computing community uses, including Autoencoders, ResNets, and Transformers. Given our extensible infrastructure, the users can also extend and build their own bundles by adding to the Python front-end and C runtime.
Figure 6 demonstrates the layers in ResNet-50 grouped into a list of bundles. Some bundles only have a Conv2D layer, while others include activations, pooling, flattening, and skip connections. Our intermediate representation realizes this by optionally enabling the different layers within a bundle. The Python front-end handles the transformation, verifies the intermediate outputs by comparing with qkeras results, statically partitions them to be run on the CGRA and the CPU, and generates C firmware to execute the bundles in runtime.
Convolutional and dense layers and matrix multiplications are repetitive but computationally heavy but tensor operations that require hundreds or thousands of MAC (multiply-accumulate) to compute a single summed result. These are best suited to be accelerated in our dedicated, parameterized CGRA with $R$ rows and $C$ columns of processing elements that compute $R{\times}C(=192,768,1024)$ summed values in parallel over hundreds or thousands of clocks before storing them in on-chip memory.
Pixel-wise operations like pooling are lightweight but complex since countless edge cases need to be handled to replicate the behavior of ML frameworks like Keras. Therefore, these are executed in software on the CPU that controls the CGRA, while the engine computes the next set of summed pixels. This way, further operations can be easily added to the firmware without rewriting and verifying the hardware.
For more elaborate operations such as the attention mechanism, we provide classes that are internally built using multiple bundles and decomposed as such during the runtime. This helps the user build models such as transformers with ease.
Figure 6: Bundle: A given DNN is decomposed into a list of bundles, where a bundle is a group of layers that can be deterministically executed by the system generated by CGRA4ML. The CGRA accelerates the simple but compute-heavy operations, while the CPU executes the complex but lightweight pixel-wise operations. This flexibility allows us to add new, complex operations easily.
V-B CGRA4ML Software
The Python library is built upon qkeras for building and training quantized models. qkeras was chosen because the scientific computing community widely uses it to build and implement models on FPGAs. A bundle is implemented in Python as a subclass of a qkeras Layer, such that a list of bundles form a QModel that can be natively trained using the optimization algorithms available in qkeras [44]. qkeras supports fixed point, binary, and ternary quantization techniques, with various scaling and rounding options, such as rounding to the nearest even and stochastic rounding, to achieve numerical stability and to maximize accuracy in the smallest number of bits. Qkeras also provides quantized and hardware-friendly approximations to activation functions such as ReLu, Softmax, Sigmoid, and Tanh. During backpropagation, qkeras uses a straight-through estimator [45] to maintain stability.
After the quantization-aware training, our framework decodes the graph structure of the given neural network. The directed acyclic graph is traversed depth-first to identify dependencies between layers such output tensors and skip connections. The runtime buffer sizes for these tensors are determined based on the tensor shapes, and their lifetimes are calculated from the graph traversal to identify at which point they should be freed during the runtime. Then, each tensor in the generated runtime is statically assigned a memory address, freeing and reusing the buffers after their lifetimes. This compile-time static allocation makes the C runtime deterministic, avoiding memory leaks.
V-C CGRA4ML Firmware
CGRA4ML provides an optimized runtime written in generic C that can be compiled to ARM, RISC-V, or other targets. The Python front-end writes a config.h file that contains the model definition as a list of bundles with runtime parameters. The C runtime, compiled by the user with the model-specific config.h, handles the engine, reading inputs and writing outputs to pre-determined buffers. This allows the user to add further co-processors and Ethernet/PCIe pipelines by simply moving the data to and from these buffers.
Our firmware is built in three layers, to be modular and extensible to new architectures & embedded systems:
  • Model specification - exported from Python API
  • Runtime - independent of the architecture & model
  • Architecture specific functions
The runtime is written as a set of C functions that can be compiled and executed for any architecture (x86, ARM, RISC-V), and can be interfaced with the testbench as DPI-C functions for any simulator: Vivado XSim, Synopsys VCS, Verilator (C++). The firmware runs in parallel to the CGRA, applying pixel-wise operations to the tensor computed by the CGRA in the previous iteration. Two banks of on-chip memory are used in a ping-pong configuration. Data synchronization between the host CPU and the CGRA is achieved by acquiring an releasing the lock registers implemented in the AXI-Lite configuration register bank.
V-D CGRA4ML Verification
Verification is an often overlooked aspect of research designs, which results in significant friction when transitioning from FPGA to ASIC flows. When users implement sophisticated models and developers add new features, it is important to ensure they will work as intended on the final embedded system before investing the effort to implement them. Our verification infrastructure has been tested on Synopsys VCS, AMD-Xilinx Xsim, and Verilator (open source).
CGRA4ML offers a sophisticated verification suite that is integrated across hardware, software, and firmware. Our randomized, transactional testbench suite is built in SystemVerilog and interfaces with the IP’s AXI-Manager ports. The SystemVerilog Direct Programming Interface for C (DPI-C) is utilized in a novel manner to interface the runtime firmware functions with the SV testbench suite to achieve holistic verification.
First, a randomized tensor is passed through the user’s deep neural network model to capture intermediate and final output tensors in the software frontend. These are sliced, reshaped, and tiled to produce the expected inputs and outputs from the dataflow. During simulation, the data memory resides on the C side and is accessed from the SV side through custom functions. When the simulation begins, a function from the C runtime is called through DPI-C to load the model-specific configuration into the AXI-Lite register bank. As the simulation progresses, the AXI interfaces of the SV testbench read the data memory from the C side and feed the tensors into the RTL design. The outputs from the design are similarly stored into the memory. This process is randomized, such that the valid & ready signals of the address, data and response channels of the AXI interfaces are randomly toggled at the probability set by the user. This randomization emulates the behavior of a congested bus interface, stress-testing the system under realistic conditions and revealing bugs that would otherwise be unidentifiable. The intermediate and final results from the simulation are dumped into files, which the Python API then tests against the expected results.
Our comprehensive verification suite ensures that the model built and trained by the user works as it should on the RTL design and the firmware of the entire embedded system, before even starting the implementation process. After implementation, the same suite can be used on the gate level, time annotated netlist. This allows the developer and the user to add functionality to hardware and the runtime firmware and test it on their desktop before testing it on an FPGA, allowing agile development cycles. A set of GitHub actions running in the cloud verify the entire infrastructure using Verilator for multiple random examples after every update, enabling continuous integration and development (CI/CD), as shown in Figure 5.
V-E FPGA & ASIC Toolflows
CGRA4ML is built to generate synthesizable SystemVerilog RTL hardware to support both FPGA and ASIC implementation flows while being agnostic to vendor-specific tools and devices. Our hardware modules feature open AXI, AXI-Stream, and AXI-Lite interfaces to ensure compatibility with a wide variety of SoCs and other hardware IPs. We utilize open-source AXI and AXI-Stream cores [46, 47] in our SoC generation, modified to support ASIC flow in addition to FPGA. All these modules are wrapped into a single IP that can be easily instantiated and connected to any system with AXI ports.
V-E1 FPGA Toolflow
Our design decisions ensure that integrating the generated IP into an FPGA system is fairly straightforward. We provide modular TCL scripts to enable such integration, which automates the following process. On AMD-Xilinx FPGAs, the ZYNQ processing system is first configured to have three full AXI Subordinate ports with the maximum width (128-bit on ZCU104) for the highest throughput and one AXI-Lite Manager port for configuration. Their address mappings are updated, and the IP ports are connected to them. The TCL scripts then verify the connections, synthesize, place & route the design, exporting bitstream and generating reports.
In addition to this standard flow, the standardized AXI interfaces of CGRA4ML allow the user to integrate the generated IP into existing signal processing dataflow systems easily. Our scripts for AMD-Xilinx Vitis then create an application project, import the bitstream and C firmware, set compiler optimizations and allow the user to execute the system as baremetal.
PetaLinux, the AMD-specific embedded Linux distribution, is widely used for higher-level applications with ZYNQ systems. When executed on PetaLinux, our C firmware manages the physical to virtual address mapping to provide seamless operation with an intuitive execution API, the same as baremetal. Our C firmware is also wrapped in Python using ctypes, allowing users to pass and receive data as Numpy arrays in a Python environment. This PYNQ API allows the user to be productive during prototyping, similar to the PYNQ API offered by hls4ml. However, as the user moves towards production, our framework also allows layers of the same API to be extended into an optimized production-ready C runtime.
V-E2 ASIC Toolflow
ASIC design is a process that often takes several months to years. To obtain a first-order approximation of power-performance-area (PPA) analysis of a given CGRA configuration, we provide a basic set of TCL scripts for Cadence Genus & Synopsys DesignCompiler for synthesis and Cadence Innovus for place & route. Users can link their PDKs, set a few basic parameters, and run the scripts to generate GDSII files and reports. More experienced ASIC designers can further tweak the scripts for a more optimized design.
We have partnered with ARM to tape out an SoC with the IP generated by CGRA4ML integrated into NanoSoc [48], an open-source SoC platform around an ARM-M0 processor. The platform features APB ports for configuration and three AHB Subordinate ports for data movement in its expansion region. To make our design compatible with their system, we add AXI to AHB bridges around our design and a custom APB Subordinate port in place of the AXI-Lite port.
VI Results
In this section, we present the results of experiments comparing CGRA4ML to hls4ml on ZCU104 evaluation board with Zynq UltraScale+ XCZU7EV-FFVC1156-2-E MPSoC FPGA. We also demonstrate CGRA4ML on deeper models such as ResNet-50 and PointNet, which are too large to be implemented using hls4ml. Our ASIC results show the power and area efficiency of the RTL-based CGRA implemented using the ASIC flow provided with CGRA4ML.
Fig. 7 demonstrates the resource and power consumption for different PE arrays on the FPGA, with rows=8. Since larger arrays are more efficient in terms of GOPs/W, users should strive to implement the largest possible array within the available resources.
[…]
Figure 7: Scaling of hardware resource and power consumption on a Xilinx ZCU104 FPGA, for different bitwidths and PE columns, with PE rows = 8
Figure 8: Performance analysis of ResNet-50 on different CGRA array sizes. Performance efficiency measures how well each layer is dynamically mapped over statically configured hardware. While larger arrays reduce off-chip data movement due to high data reuse, they have lower utilization for smaller layers. The unified dataflow of the array (Sec. IV-B) maximizes the utilization to near 100% and minimizes data movement for most layers.
Model ResNet-50 PointNet
Bits 4 4
PEs (7,96) (32,32)
Frequency (MHz) 250 250
FFs 101706 69277
LUTs 82200 100076
BRAMs 6 4.5
Static Power (W) 0.700 0.700
Dynamic Power (W) 3.847 3.840
Total Power (W) 4.547 4.540
GOPs/W 37.3 56.8
TABLE III: Implementation of ResNet-50 and Pointnet on ZCU104 FPGA
VI-A ResNet-50
ResNet-50 [49] is a deep convolutional neural network with 5.36 billion operations in 53 convolutional layers, one dense layer, two pooling layers, and 17 skip connections. Its input is a $224{\times}224{\times}3$ tensor, and outputs are 1000 classes, and it is run on an array of $7{\times}96{=}672$ PEs.
[…]
Framework CGRA4ML hls4ml
Reuse factor - - 1 16 64
PEs (8,12) (16,32) - - -
Frequency (MHz) 250 250 250 250 250
FFs 32451 64066 21428 20991 21335
LUTs 17959 43320 11810 11986 11963
BRAMs 4.5 4.5 5 5 7
Static Power (W) 0.694 0.697 0.693 0.693 0.693
Dynamic Power (W) 3.078 3.450 2.948 2.886 2.820
Total Power (W) 3.772 4.147 3.642 3.579 3.513
TABLE IV: Comparing Jet-tagger model on CGRA4ML with different array sizes and on hls4ml with varying reuse factors
VI-B PointNet & Jet Tagger
PointNet [50] is a model developed for particle classification as a stack of three 1D convolutions, three dense layers, and two pooling layers. An input size of 10.6K requires 1.18 billion operations for a batch size of 16. We demonstrate this on an array of size $16{\times}32{=}512$. Jet Tagger is a small model with 117K operations in 4 dense layers, widely used for benchmarking. Table IV compares the resource and power consumption of hls4ml and CGRA4ML for the Jet tagger model.
VI-C ASIC Implementation
Technology 7nm 65nm
PEs 768 192 192
Frequency (MHz) 1000 1000 250
Area (mm2) 0.164 0.049 0.47
Static power (mW) 23.62 7.38 54.04
Dynamic power (mW) 381.1 119.9 26.80
Total power (mW) 404.7 127.2 80.84
GOPs/W 1898 1509 594
TABLE V: ASIC results from Cadence toolflow with 7nm and 65nm TSMC LP CMOS PDKs for CGRA configurations with $8{\times}24$ and $8{\times}96$ rows and columns generated using CGRA4ML
Figure 9: CGRA of $R{\times}C{=}8{\times}24$ configuration generated by CGRA4ML, synthesized with Cadence Genus and placed & routed with Cadence Innovus using TSMC 65nm LP CMOS PDK, being taped out with NanoSoC platform.
As described in Sec. V-E2, we are collaborating with Arm Research to tape out an SoC built through CGRA4ML. We utilize NanoSoC[48], an open-source, ARM-based SoC generation platform built in collaboration with ARM Research. Our CGRA is integrated with the ARM processing system through the DMA360 open-source IP developed and provided in the NanoSoC framework that connects three AXI-Stream interfaces to the AHB interfaces of the NanoSoC interconnect. The ASIC generation toolflow of CGRA4ML is demonstrated through the results presented in Table V. The CGRA from CGRA4ML delivers high performance in a few milliwatts, making it suitable for scientific computing applications at the edge. Our $8{\times}24$ variant of the CGRA with 192 PEs, generated by CGRA4ML, is being taped out into an SoC with the TSMC 65nm LP CMOS PDK.
VII Conclusion
In this paper, we introduced CGRA4ML, a framework for implementing large-scale neural networks in scientific edge computing applications. CGRA4ML successfully bridges the limitations of existing solutions by facilitating off-chip data storage and supporting a diverse array of neural network architectures, including but not limited to Autoencoders, ResNets, and Transformers. Our framework intends to complement hls4ml, a tool widely used by the scientific computing community by handling larger and more complex models that are crucial for advanced scientific computing tasks.
The ability of CGRA4ML to generate RTL designs and runtime firmware, and support end-to-end, vendor-agnostic ASIC and FPGA design flows is a notable advancement, providing a robust solution for researchers and developers aiming to leverage the speed and efficiency of hardware acceleration without the traditional complexity. We demonstrate the effectiveness of our framework by comparing it with hls4ml.
Our contributions are not only technical but also extend to the ease of use and extensibility, thanks to CGRA4ML’s modular and minimal infrastructure. Our Python API, along with its comprehensive verification system, ensures that scientific researchers can focus more on innovation and less on the intricacies of hardware design.
By detailing the framework’s architecture, implementation, and potential for future enhancements, this paper provides a valuable resource for the ongoing development of edge computing technologies. We believe CGRA4ML will play a crucial role in advancing the capabilities of scientific edge computing, paving the way for more sophisticated and real-time data processing at the edge.