### v3

# cgra4ml: A Hardware/Software Framework to Implement Neural Networks for Scientific Edge Computing

Abstract.
The scientific community increasingly relies on machine learning (ML) for near-sensor processing, leveraging its strengths in tasks such as pattern recognition, anomaly detection, and real-time decision-making. These deployments demand accelerators that combine extremely high performance with programmability, ease of integration, and straightforward verification. We present cgra4ml, an open-source, modular framework that generates parameterizable CGRA accelerators in synthesizable SystemVerilog RTL, tailored to common ML compute patterns found in scientific applications. The framework supports seamless system integration through AXI-compliant interfaces and open-source DMA components, and it includes automatic firmware generation for programming the accelerator. A comprehensive verification suite and a runtime firmware stack further support deployment across diverse SoC platforms. cgra4ml provides a modular, full-stack infrastructure, including a Python API, SystemVerilog hardware, TCL toolflows, and a C runtime, which facilitates easy integration and experimentation, allowing scientists to focus on innovation rather than dealing with the intricacies of hardware design and optimization. We demonstrate the effectiveness of cgra4ml to implement common scientific edge neural networks using ASIC and FPGA design flows.
FPGA, reconfigurable computing, edge computing, machine learning, CGRA
journal: TRETSccs: Hardware Reconfigurable logic applicationsccs: Computing methodologies Machine learning
1. Introduction
Modern scientific discovery increasingly relies on machine learning to process vast volumes of data at the edge, where sensors capture events in real-time and decisions must be made within microseconds (Duarte et al., 2018). From high-energy physics experiments and remote sensing arrays to in-situ biomedical diagnostics and fusion plasma control, these scientific applications demand neural network inference on custom hardware architectures to achieve the required inference speed and energy efficiency with extremely high throughput, low latencies, and tight energy budgets (Coelho et al., 2021). Such custom hardware designs must be closely integrated with host microprocessors and memory subsystems to achieve the highest performance. At the same time, these workloads increasingly utilize complex, evolving neural network models, necessitating a degree of programmability and architectural flexibility. Previous scientific edge computing frameworks such as hls4ml (FastML Team, 2024) and finn (Umuroglu et al., 2017) have enabled small neural networks on FPGAs through high-level synthesis. However, they struggle to scale to deeper models, offer limited flexibility in hardware reuse, and lack production-ready verification and integration flows. Therefore, the scientific computing community is in need of a framework that supports larger models with a user-friendly interface, support for various quantization levels, modular hardware architecture, production-ready firmware generation, complete system integration, and verification.
To address these requirements, we present cgra4ml, an open-source framework for deploying modern neural networks in scientific edge computing environments. cgra4ml generates parametrizable, reusable Coarse-Grained Reconfigurable Array (CGRA) hardware engines and associated runtime systems, enabling scalable inference of complex models across FPGAs and ASICs. The framework delivers a holistic solution, comprising a Python-based frontend for quantized model specification, a SystemVerilog RTL backend for the CGRA, a modular C runtime firmware, and an integrated verification suite.
cgra4ml is targeted to two user groups: (1) Scientists, embedded / IoT researchers who want to build and implement modern neural networks that cannot be deployed using current popular frameworks for scientific computing due to their size and complexity. (2) Hardware system designers who want to iteratively test new optimizations such as new datatypes, dataflow, and compression techniques.
The major contributions of cgra4ml include:
  • A hardware + software + firmware co-design framework to specify, train, quantize, and implement diverse neural networks on FPGA and ASIC as CGRA-powered SoCs.
  • An open-source CGRA architecture with dynamic reconfiguration, unified dataflow, and an efficient PE design to maximize hardware utilization for modern neural networks.
  • RTL-based high-performance, modular SoC design with open-source DMAs and AXI-compliant interfaces to support vendor-agnostic FPGA and ASIC design flows.
  • Automated firmware generation enabling hardware/software partitioning that easily integrates the CGRA with on-chip processors.
  • End-to-end implementation and test on AMD Xilinx FPGAs and ASIC flow using Cadence tools to be taped out as an SoC with ARM-based NanoSoC platform.
  • Extensive verification support to test the neural network model against the generated RTL design and C runtime through randomized transactional testbenches.
We evaluate cgra4ml on representative scientific workloads, including ResNet-50 for high-resolution image classification and PointNet for particle cloud inference. Results on AMD-Xilinx FPGAs and Cadence ASIC flows demonstrate that cgra4ml achieves competitive performance-per-watt, extensibility, and scalability, making it a robust and practical solution for real-time scientific inference at the edge.
2. Background and Related Work
This section outlines the requirements for a hardware-accelerated system development in scientific computing. We first introduce the need for neural networks in scientific computing and practical use cases. We then discuss two ways of implementing such networks: a layer-by-layer pipelined implementation where each layer becomes a pipeline stage in hardware and the accelerator-based implementation where all layers reuse the same reconfigurable hardware. hls4ml (FastML Team, 2024) and finn (Umuroglu et al., 2017) take the first approach, while several accelerators (Juracy et al., 2023) (Dhilleswararao et al., 2022) and frameworks such as Apache TVM with VTA accelerator (Moreau et al., 2018) take the second. We then outline the uniqueness and utility of a CGRA-type accelerator architecture to solve this problem and discuss related research in this area. Consequently, we summarize the needs of the scientific computing community that are not addressed by other generic frameworks as motivation for cgra4ml.
Figure 1. Data processing pipeline at Large Hadron Collider, CERN. The detectors at LHC produce data at a rate of 320 Tb/s during collisions. Due to the infeasibility of transferring data at such a high rate, the events are filtered through a multi-step pipeline. Small neural networks like autoencoders are implemented using hls4ml on radiation-hardened ASICs. More sophisticated models are implemented on FPGAs (L1) and servers (L2) to filter the data fully. With cgra4ml enabling larger models at low power, more processing can be moved to the edge (towards the left), resulting in more robust filtering.
Data processing pipeline at Large Hadron Collider, CERN.
2.1. Neural Networks for Scientific Computing
Machine learning is fueling scientific discoveries in particle physics, materials, dark matter, cosmology, nuclear physics, biomedical engineering, and health monitoring (Deiana et al., 2022). Specific examples include bio-signal classification (Ahmad and Zafar, 2024), tracking magnetohydrodynamic (MHD) instabilities in fusion reactors (Wei et al., 2023), reinforcement learning for accelerator beam control (Duarte et al., 2022), pulse detection for anti-neutrino detection (Axani et al., 2024), and accelerator controls for beam loss deblending (Arnold et al., 2023). Scientific computing increasingly relies on neural networks to process data with requirements for extremely low latencies and high throughputs.
The Large Hadron Collider (LHC) has $\mathcal{O}(100M)$ individual sensors generating data for each of the 40 million proton beam collisions per second (Qu et al., 2022). LHC trigger systems filter events with sub-microsecond latency and 40 MHz throughput requirements using FPGAs (Duarte et al., 2018) as shown in Fig. 1. In each level, data is compressed, analyzed, and filtered. While traditional algorithms have been used for data filtering in the past, neural networks accelerated in FPGA and ASICs are replacing them. For example, neural network-based autoencoders are implemented as ASICs in the L0 trigger to compress and move data out of the sensors (Di Guglielmo et al., 2021). L1 uses more complex networks accelerated on FPGAs, and L2 triggers and offline reconstruction use even bigger models executed on GPUs. User-friendly frameworks for hardware implementation have fueled the wide adoption of small neural networks in lower-level triggers. hls4ml is one of the more popular among scientific edge accelerator frameworks.
Large models, such as Particle Transformer (Qu et al., 2022) for particle jet classification, are being explored by the physics community. However, for low-latency FPGA trigger use, they have not yet shown compelling gains, so practical deployments remain GPU-based. This is partly a chicken-and-egg issue. Since hls4ml implementation of models requires more and more resources with increasing model size, almost all the effort from the scientific computing community goes into making their models smaller to fit within an FPGA. Multi-FPGA routes, such as AIgean (Tarafdar et al., 2022), point to the possibility of using larger models but are comparatively cumbersome for trigger pipelines.
2.2. Dataflow-style NN Implementation: hls4ml, finn
Figure 2. Layer-by-layer dataflow-style implementation of neural networks using hls4ml. hls4ml offers multiple backends, one per family of devices. Each backend is a collection of layers defined as templated HLS. Each neural network layer is implemented as a separate datapath. Xilinx DMA IPs are connected to the HLS4ML design to move data in and out. finn has a similar dataflow-style implementation. In contrast, cgra4ml reuses the same CGRA multiple times to process a layer.
Layer-by-layer spatial implementation of neural networks using hls4ml
hls4ml is a Python library developed low-latency, high-throughput edge machine learning inference (FastML Team, 2024). hls4ml primarily supports models built and trained using QKeras, a library from Google for low-precision machine learning (QKeras Team, 2024). hls4ml converts each layer of the model into a customized high-level synthesis (HLS) hardware block and connects them together to form an IP for the entire model (see Figure 2). The user controls reuse factor per layer, which defines the number of times a multiplier/accumulator is reused. The reuse factor provides a way to control a layer’s resource usage and performance. Balancing performance and resource utilization is highly dependent on the specific target FPGA and involves iterative manual tuning of reuse factors on a per-layer basis. This optimization effort is non-trivial and takes a significant time and effort.
Although the layer-by-layer approach may offer advantages in latency minimization, the architecture suffers from scaling issues as the size of its target NN increases. Because each layer is implemented with a unique datapath, larger NNs with tens to hundreds of layers quickly become too complex to implement on even the largest of hardware devices.
As the name implies, hls4ml requires a high-level synthesis (HLS) tool flow to generate the hardware accelerator. hls4ml has different HLS backends (AMD Xilinx FPGA, Altera FPGA, Siemens Catapult, etc.) with different levels of support. The most common target is AMD Xilinx Vitis/Vivado HLS, though other backends exist in various forms (Altera HLS Compiler, Siemens Catapult HLS, etc.). The dependence on HLS restricts the user to the supported vendors and requires HLS tool-specific reimplementation of each hardware layer, complicating the maintenance and verification. hls4ml’s support for ASIC implementation is primarily through Catapult HLS, which currently only supports a limited subset of layers.
A common hls4ml design flow requires scientists to generate an IP for the neural network, and then engineers write firmware, perform hardware-software integration, verification, and debugging, which is an onerous task. There are efforts to streamline this process for the scientists. For example, hls4ml provides a PYNQ overlay, a Python-based runtime for AMD Xilinx Zynq FPGAs. This workflow allows the physicists to focus on building and training low-precision models for their custom applications without spending months on hardware design, implementation, and verification. Yet, by and large, the ML IP core integration for production-ready designs remains a separate process from hls4ml.
finn (Umuroglu et al., 2017) is another popular open-source FPGA accelerator framework that uses a similar methodology. finn solely targets FPGAs though its general methodology could be ported for other hardware backends. finn shares many of the benefits and drawbacks of hls4ml, while being especially optimized for extremely low bitwidths, such as binary neural networks (Borras et al., 2022). The layer-by-layer approach is ideal for minimizing latency, since it eliminates the need to transfer partial outputs to/from the off-chip memory. It also makes the NN implementation extremely modular, allowing any layer to be added in any order. Yet, finn is primarily focused on generating an IP core that must be separately integrated into a larger system.
2.3. DNN Accelerators
Figure 3. Positioning CGRA4ML in the space of ML-to-FPGA frameworks from the perspective of the scientific computing community. hls4ml (FastML Team, 2024) is the most popular tool in the scientific computing community, as it is user-friendly and supports the dataflow-style implementation of models of varying bitwidths and layers to an extent. finn (Umuroglu et al., 2017) implements models in a similar dataflow fashion, but excels at very low bitwidths. Traditional AI accelerators are suitable for large models with 8+ bit quantization. cgra4ml aims to fill the gap by making models with sub 8-bit quantization that are too big to implement with hls4ml and finn.
Layer-by-layer spatial implementation of neural networks using hls4ml
Hundreds of accelerator architectures have been designed for deep neural network inference in the past decade (Peccerillo et al., 2022; Juracy et al., 2023; Dhilleswararao et al., 2022). This has been the default approach for the implementation of DNNs on hardware. These accelerators can be categorized by their data reuse pattern as weight, input, output and row stationary and by their runtime flexibility into systolic arrays, CGRAs, and microcode processors.
Caffeine (Zhang et al., 2019) offers an HLS systolic array to process CNNs from the Caffe ML framework using weight major and input major mappings. Eyeriss (Chen et al., 2017) is an energy-efficient ASIC designed to accelerate CNNs using a 2D array of 168 fairly complex processing elements, each with a 16-bit multiply-accumulate, scratchpad, and a controller. ShiDianNao (Du et al., 2015) uses a 2D mesh of functional units optimized towards the 2D feature maps of convolutional layers. Each processing element executes multiplications, additions, and comparisons using 16-bit fixed point arithmetic. The kernel elements are shifted right to left and up to down and accumulate locally.
While these architectures demonstrate strong performance and efficient data reuse, they often lack the broader system-level features required for scientific edge applications. Many of these accelerators are optimized for fixed model types, offering limited programmability and making it challenging to re-target them for newer or more diverse neural networks. Integration into embedded systems is frequently an afterthought, as they rarely include clean, general-purpose interfaces for control or data exchange with a microprocessor, which complicates deployment in real-time, heterogeneous environments. Additionally, verification support is typically constrained to smoke tests, presenting challenges for researchers and engineers who need robust testing and validation pipelines for use in high-stakes scientific contexts. Our framework addresses these gaps by combining performance-efficient hardware with runtime programmability, a microprocessor-friendly interface for seamless system integration, and a comprehensive verification suite tailored to the demanding requirements of scientific edge computing.
2.4. CGRA
Coarse-Grained Reconfigurable Arrays (CGRA) is a type of architecture that has seen recent resurgence in industry and academic designs. Commercial CGRAs include Samsung Reconfigurable Processor (Kim et al., 2012) and the Renesas STP Reconfigurable Processor (Corporation, ). CGRA-ME (Chin et al., 2017) is an exemplary academic framework that supports a subset of CGRA designs.
CGRAs were originally envisioned as coarse-grained FPGAs whose programmable elements work at the word-level instead of FPGA bit-level programming (Liu et al., 2019). Early CGRAs were classified based on their integration into the processor. Tightly coupled CGRAs integrate into a processor data path and are executed as a custom instruction, e.g., Chess (Marshall et al., 1999), MATRIX (Mirsky and DeHon, 1996), and DySer (Govindaraju et al., 2012). Loosely coupled CGRAs act more as an accelerator that executes alongside the processor, executing in tandem with the processor and communicating via on-chip interconnect. Examples of loosely-coupled CGRAs include PipeRench (Goldstein et al., 2000), MorphoSys (Singh et al., 2000), CHARM (Cong et al., 2012), and FPCA (Cong et al., 2014). cgra4ml can be characterized as a loosely coupled CGRA as it implements sub-tasks of the neural network alongside a CPU with a focus on enabling hardware/software partitioning across the processor and processor.
Modern neural network accelerators often exhibit architectural similarities to CGRAs, particularly in their use of spatial dataflows, on-chip communication networks, and pipelined compute units. Like CGRAs, many NN accelerators rely on configurable interconnects and local memory structures to maximize data reuse and throughput. However, neural network accelerators are typically more narrowly optimized for specific computational patterns and may hardwire aspects of control or scheduling for performance. In contrast, CGRAs are generally more flexible—supporting a broader range of computational kernels and allowing dynamic reprogramming of data paths at runtime. With cgra4ml, we aim to navigate the design space between these two classes: delivering an architecture that captures the efficiency and specialization of NN accelerators while retaining the programmability and reuse benefits of CGRAs.
In the recent literature, OpenCGRA (Tan et al., 2020) introduced a unified, open-source stack to model, test, simulate, and characterize CGRAs, with LLVM compiler passes and PyMTL-based Verilog generation, such that designers can quickly explore heterogeneous tiles and full EDA flows in one place. AURORA (Tan et al., 2021) then automated CGRA co-design: starting from a generic architecture template, it jointly explores composition of functional units, memory/interconnect, and loop optimizations using an architecture-aware simulated annealing search. VecPAC (Tan et al., 2023) proposes a hybrid CGRA supporting different datatypes with scalar tiles plus configurable vector tiles. PICACHU (Qin et al., 2025) introduces a plug-in CGRA for LLM nonlinear ops (GELU/Softmax/LayerNorm, etc.) with domain-specialized PEs, precision-aware design, and seamless integration with a systolic array. ML-CGRA (Luo et al., 2025) is an MLIR-based end-to-end framework that keeps ML-level semantics and adds passes lower whole models onto CGRAs. AHA introduces an agile CGRA-compiler co-design flow where three domain specific languages: PEak for PEs, Lake for memories, and Canal for interconnects, auto-generate both RTL and compiler collateral from a single source of truth, so the compiler updates automatically as hardware evolves. While such CGRA frameworks primarily target architecture exploration and compiler lowering, cgra4ml’s use case is complementary and deployment-oriented. It delivers vendor-agnostic SystemVerilog RTL and a production C runtime verified in-the-loop via DPI-C, packaged as AXI IP with DMA that can be integrated with SoC platforms such as Zynq, NanoSoC and Ibex, enabling rapid, reproducible deployment in heterogeneous SoCs.
Table 1. Comparing cgra4ml with existing ML-to-FPGA/ASIC frameworks. As a deployment-oriented framework, cgra4ml’s unique feature is its comprehensive verification approach, which tests the user’s model and the production runtime on the generated hardware via DPI-C with randomized SoC-level congestion emulation. In addition, we provide a superset of features from other frameworks in one place for the scientific computing community
This work hls4ml (FastML Team, 2024) VTA (Moreau et al., 2018) FINN (Umuroglu et al., 2017) Vitis AI (AMD-Xilinx, 2024)
Open-source
Sub-8-bit quantization support
Larger neural networks support
Xilinx FPGA support
Intel FPGA support $\sim$
ASIC Implementation $\sim$
Basic verification
Generate production-ready runtime
Verification with runtime
DLP (Mathworks, 2024) OpenVino (Intel, 2024) Caffeine (Zhang et al., 2019) CGRA-ME (Chin et al., 2017) FlexCNN (Basalama et al., 2023)
Open-source
Sub-8-bit quantization support
Larger neural networks support
Xilinx FPGA support
Intel FPGA support
ASIC Implementation
Basic verification
Generate production-ready runtime
Verification with runtime
2.5. ML to FPGA/ASIC Frontend Frameworks
There are many end-to-end frameworks available for hardware implementation of neural networks (Rahimifar et al., 2022). DNNBuilder (Zhang et al., 2018) and finn (Umuroglu et al., 2017), implement a given model as a pipeline of layers, similar to hls4ml. DNNBuilder allows the users to customize two kinds of reuse factors: channel and kernel. finn takes in an ONNX model, possibly exported from Brevitas (Pappalardo, 2024), to generate Vivado HLS IP. This can be verified in simulation before implementing on AMD Xilinx FPGAs. FINN provides a PYNQ driver for prototyping.
AMD Xilinx Vitis AI (AMD-Xilinx, 2024) is a closed-source library that implements a Deep Learning Processing Unit, an 8-bit micro-coded processor to process neural networks optimized through their stack on AMD Xilinx FPGAs. OpenVino is a similar stack for Intel FPGAs. Apache TVM, an open-source framework for embedded AI, implements a Versatile Tensor Accelerator (VTA) (Moreau et al., 2018) as a GEMM processor using Xilinx HLS. LeFlow (Noronha et al., 2018) emits HLS code, which has similar advantages and disadvantages as hls4ml.
Recent compiler stacks such as PyTorch Dynamo (Team, 2025b) and OpenAI Triton (Team, 2025a) focus on automatic kernel fusion and code generation for GPUs/CPUs (with early explorations for other backends). Our contribution is orthogonal: cgra4ml targets a CGRA-style spatial accelerator with unified dataflow and a firmware/runtime that binds those schedules to a memory-mapped, DMA-driven engine on FPGA/ASIC. In principle, TorchInductor/Triton-style schedules could be lowered to our ISA/dataflow with a translation layer; we leave such integration to future work.
While each of these frameworks targets different kinds of users, the limitations in their approaches make them incompatible with the requirements of the scientific computing community, as listed in Table 1. The popularity of hls4ml and its features: arbitrary quantization, Xilinx and Intel FPGA backends, and support for limited verification, demonstrate the unique needs of the scientific edge community. The models used in scientific applications require quantization to arbitrary bit-widths, which is made possible by qkeras, and not supported by Vitis AI, OpenVino, or CGRA-ME, among others. In addition, their workflow involves implementing their models on FPGAs and later moving to ASIC designs, which is not possible with Vitis AI, Apache VTA, and OpenVino since their backends are implemented in vendor-specific HLS.
2.6. Motivation for cgra4ml
Modern scientific applications increasingly rely on neural networks to process data rapidly and efficiently at the edge, imposing stringent requirements on throughput, latency, and power efficiency. hls4ml is a popular framework among scientists—deployed in custom ASICs for L1 triggers and on FPGAs in L2 triggers at the LHC, among other uses—because it offers an accessible, dataflow-style, layer-by-layer implementation for small networks and a range of precisions. However, this approach scales poorly beyond a few layers: on-chip buffering and per-layer specialization grow quickly, making larger models difficult to fit on FPGAs and small ASICs. Reusable accelerators amortize resources across layers and are therefore more suitable for larger networks, yet most available designs provide limited, non–user-friendly front ends and often lack features needed by the scientific computing community, including sub–8-bit quantization, multiple backends (FPGA and ASIC), end-to-end verification, and runtime generation. Moreover, hls4ml and other HLS-based DNN frameworks are tied to vendor-specific toolchains, which constrains backend portability and demands substantial engineering to build complete systems with a host CPU, DMA subsystems, drivers, and production firmware.
As summarized in Fig. 3, hls4ml remains the community’s go-to for dataflow implementations over varying bitwidths and shallow-to-moderate depths, while finn likewise follows a dataflow paradigm and excels at ultra–low-precision designs. Traditional AI accelerators, in contrast, target large models with $\geq$8-bit quantization. cgra4ml is designed to fill the gap between these regimes by enabling models with sub–8-bit quantization that are too large to realize with hls4ml or finn, while providing multiple backends, end-to-end verification, and automatic runtime/driver generation to streamline full-system integration.
3. cgra4ml Overview
The needs of the scientific computing community for a framework that can implement high-performance and programmable accelerators using the familiar hls4ml-style interface, with the generation of cross-platform hardware and production-ready firmware, were identified in Section 2. Often models exceed on-chip memory capacities, demanding efficient off-chip data handling and more flexible hardware configurations. And typically users would default to not using larger models due to the inability for hls4ml to implement them. To address these limitations, we introduce cgra4ml, an open-source, modular framework explicitly designed for the scientific edge computing community that can better facilitate reuse across layers, implement larger models, and ease integration into larger systems. cgra4ml leverages the strengths of hls4ml: ease of use, quantization-aware training, and rapid hardware accelerated deployment, while substantially extending capabilities in several key areas.
Firstly, cgra4ml supports large-scale neural network implementations by employing a parametrizable CGRA, enabling efficient spatial reuse of processing elements and optimized off-chip data movements. cgra4ml outputs SystemVerilog RTL eliminating the need for HLS tools required by other frameworks. This RTL-centric approach streamlines FPGA implementation and facilitates seamless ASIC integration, thereby providing flexibility across different hardware targets without extensive code modifications. Furthermore, cgra4ml includes an extensive verification framework integrated within its workflow, significantly reducing the time spent in debugging and validation processes, which are notoriously time-consuming in hardware design. This section describes the user workflow of cgra4ml and the Python frontend that makes our infrastructure possible.
Refer to caption cgra4ml workflow
Figure 4. cgra4ml workflow as outlined in Sec. 3. Users first build quantized neural networks and train them in a quantization-aware manner using qkeras (Sec. 4.1). Users then define a CGRA definition, which generates vendor-agnostic SystemVerilog RTL hardware specifications and TCL tool flows (Sec. 4.1). The model can then be exported to generate weights and a C runtime firmware (Sec. 4.2.2). The generated hardware IP is then verified comprehensively with the model and firmware (Sec. 10) using our randomized, transactional SystemVerilog testbench suite with DPI-C extensions. Finally, the bitstream generated from the FPGA toolflow and the C firmware can be loaded into an FPGA to be tested in seconds (Sec. 5.3). After such rapid prototyping, the same hardware design can be moved to ASIC (Sec. 5.4).
The overarching goal of cgra4ml is to empower scientists and hardware designers by abstracting away intricate hardware optimization details that slow down development for non-expert users. cgra4ml’s user workflow is intuitive, streamlined, and flexible, supporting rapid development cycles and iterative experimentation. Users define neural network architectures via a Python API leveraging qkeras for quantization-aware training, after which cgra4ml automatically generates hardware specifications, runtime firmware, and robust verification suites. This end-to-end approach enables rapid prototyping, flexible experimentation, and efficient deployment, ultimately allowing users to prioritize innovative applications over hardware intricacies. It comprises the following steps:
3.0.1. Model Definition and Training
Users begin by defining neural network models using the Python-based qkeras library, which supports quantization-aware training. This step ensures models are optimized for low-precision hardware implementations while maintaining accuracy.
3.0.2. Accelerator Configuration
After training, users specify static parameters for the CGRA architecture, including the number of processing elements (PEs), bitwidths, and memory hierarchies. Users can adjust these parameters to tune the design to within the resource constraints of the target hardware.
3.0.3. RTL Generation
Using the specified parameters, cgra4ml automatically generates vendor-agnostic SystemVerilog RTL code. This eliminates dependency on vendor-specific tools and simplifies porting between FPGA and ASIC targets.
3.0.4. Firmware Generation
cgra4ml produces optimized runtime firmware, facilitating seamless integration and operation of the generated hardware within an SoC environment. Users can quickly deploy and test their designs in real-world applications.
3.0.5. Verification
The generated hardware and firmware undergo comprehensive verification using cgra4ml ’s built-in randomized transactional testbench suite. This step ensures functional correctness and reliability prior to physical deployment.
3.0.6. FPGA/ASIC Deployment
Finally, users deploy their verified designs onto FPGA platforms for rapid prototyping and experimentation. Once validated, the same RTL design seamlessly transitions to ASIC flows.
Through this structured, hls4ml-inspired workflow, cgra4ml enables users to rapidly move from model concept to physical deployment, streamlining innovation, and accelerating scientific discovery.
4. cgra4ml: Infrastructure
To bridge the gap between high-level neural network development and hardware-accelerated edge deployment, cgra4ml provides a modular, holistic infrastructure that spans model quantization, hardware generation, firmware synthesis, and system-level verification. Central to this flow is the concept of bundles, which are modular, deterministic units of execution that encapsulate groups of DNN layers, making them suitable for hardware acceleration. These bundles allow for clean partitioning of compute workloads between a coarse-grained reconfigurable array (CGRA) and a general-purpose CPU. This design philosophy supports scalability, reuse, and extensibility, enabling developers to experiment with diverse model structures while maintaining control over system-level performance and resource utilization. The backend pipeline, which comprises SystemVerilog CGRA RTL, portable runtime firmware, and automated toolchains for FPGA and ASIC flows, ensures that scientific edge applications can transition from simulation to physical deployment with minimal effort.
4.1. Frontend Infrastructure: Model to Bundles
Figure 5. Bundle: A given DNN is decomposed into a list of bundles, where a bundle is a group of layers that can be deterministically executed by the system generated by cgra4ml. The CGRA accelerates the simple but compute-heavy operations, while the CPU executes the complex but lightweight pixel-wise operations. This flexibility allows us to add new, complex operations easily.
A given DNN is decomposed into a list of bundles
Modern neural networks have complex, seemingly arbitrarily structured graphs of computational layers. Therefore, a framework should balance the flexibility it offers to the users to the optimized backend implementation of these networks in hardware. This requires a modular and extensible intermediate representation (IR). The frontend of cgra4ml is responsible for translating a high-level neural network model into a hardware-ready specification. This process is centered around the concept of bundles, which are modular, hardware-executable units that serve as the bridge between qkeras models and cgra4ml accelerator.
Users construct and train quantized models using qkeras, a widely adopted library from Google that enables fine-grained control over quantization levels and supports straight-through estimators for backpropagation. Once trained, cgra4ml parses the model’s computational graph and decomposes it into a list of bundles. Each bundle represents a deterministic sequence of layers that can be directly mapped onto the CGRA. In the Python frontend, a bundle is implemented as a subclass of a qkeras layer, enriched with metadata and configuration parameters required for hardware mapping. In the backend firmware, each bundle corresponds to a structured set of runtime instructions, including memory access patterns, kernel parameters, and execution schedules.
A bundle acts as the core building block of the cgra4ml infrastructure and enables hardware/software partitioning. Compute-intensive operations such as convolutions and matrix multiplications are offloaded to the CGRA, while lightweight and control-heavy operations such as pooling or normalization are executed by the CPU. Convolutional and dense layers involve large numbers of multiply-accumulate (MAC) operations and benefit most from parallelization on the CGRA. The CGRA features a parameterized 2D array of processing elements ($R{\times}C$) that execute MAC operations in parallel and store results in on-chip memory. Meanwhile, CPU-executed tasks handle edge conditions and pixel-wise computations.
Figure 5 illustrates how a ResNet-50 model is decomposed into bundles, each encompassing layers such as Conv2D, activation, pooling, flattening, and skip connections. The Python frontend transforms the model, compares intermediate results against qkeras outputs, and partitions execution between the CGRA and CPU. It selectively enables layers, and passes their configuration as runtime parameters inferred from the user defined model. The generated C firmware then orchestrates bundle execution at runtime in a way that ensures functional equivalence to the Python model.
cgra4ml also empowers users to define new bundles by extending the Python frontend and adding corresponding runtime definitions in C. These composite operations are decomposed during compilation into a static runtime, allowing users to define and deploy more complex architectures such as residual networks without modifying the core hardware design. This architectural decision is made to keep the hardware complexity minimal, and to enable new operations such as activation functions when necessary. We characterize the CPU execution latency cost in Table 4 and outline our future roadmap to move the writeback stage into hardware in Sec. 6.2.
4.2. Backend Infrastructure: CGRA RTL and Firmware Generation
Figure 6. Modular and extensible infrastructure of the cgra4ml framework as described in Section 4. The Python API built around Google’s qkeras, extracts the fixed-point representations of intermediate tensors, generating the binary weights and model specification. The Python frontend also generates vendor-agnostic, synthesizable SystemVerilog RTL design of the CGRA per user specifications, optimized C runtime firmware and TCL toolflows for ASIC and FPGA implementations. Our comprehensive verification suite powered by randomized, transactional, SystemVerilog testbenches interfacing with the C firmware via DPI-C, then verifies the user’s model, the generated CGRA, and runtime together as it would run on an embedded system, emulating cases like memory congestion. The CI/CD pipeline tests the entire framework on a set of models over different CGRA configurations
Modular and extensible infrastructure of the cgra4ml framework as described
The backend of cgra4ml translates the high-level bundle specifications into hardware and runtime components, enabling deployment across FPGA and ASIC platforms. It encompasses three primary components: RTL generation, firmware synthesis, and toolchain integration for verification and implementation. This subsection outlines these steps in the backend process, while the following subsections discuss each in detail.
4.2.1. RTL Generation
cgra4ml produces synthesizable, vendor-agnostic SystemVerilog RTL for a parametrized CGRA engine. Users configure key static parameters, such as the number of processing elements, memory hierarchy depth, and data bitwidths, via the Python frontend. These parameters are embedded into the hardware description to generate an optimized datapath and control logic. The default datatype is fixed-point. However, the users may integrate a floating-point MAC by specifying its latency of the multiplier and accumulator in our parameter list. The resulting RTL supports standard AXI interfaces for seamless SoC integration. After implementation, the design must be integrated into a broader SoC. cgra4ml generates a high-performance, AXI-compliant IP powered by open-source, design tool-agnostic AXI DMAs. These modules simplify integration into custom SoCs or evaluation platforms. The firmware and hardware together ensure efficient data transfer and workload sharing between the CGRA and the host processor.
4.2.2. Firmware Generation
Alongside RTL, cgra4ml generates modular C firmware that manages CGRA execution and coordinates software-side tensor operations. This firmware is structured in three layers: (1) a model-specific configuration header exported from Python, (2) a portable runtime layer for buffer management and synchronization, and (3) architecture-specific routines for platform integration. The firmware is designed as a simple header-only library with a thin Hardware Abstraction Layer (HAL), allowing users to quickly get the design working and extend it to include co-processors and memory interfaces like Ethernet and PCIe. This modularity allows the firmware to target diverse embedded processors including ARM, RISC-V, and x86, and to be reused across different CGRA configurations. The runtime targets a stable hardware contract via a custom bank of 32-bit registers for control and status, with DMA descriptors. No changes to the unified dataflow or runtime firmware are required when targeting different architectures and FPGA/ASIC backends. Only memory macro selection and interconnect adapters vary by process and SoC fabric. We have validated the portability of our firmware on ZYNQ Programmable SoCs with FPGAs and ARM Cortex processors, on Ibex SoC with a RISC-V processor, and on x86 CPUs for verification.
4.2.3. Toolchain Integration
cgra4ml includes fully automated toolchains for FPGA and ASIC flows. For FPGAs, it generates TCL scripts that configure IPs, integrate the CGRA into SoC shells (e.g., Zynq), and synthesize/place/route the design. For ASICs, it produces scripts for Cadence Genus and Innovus, or Synopsys DC and ICC, enabling synthesis, physical design, and GDSII export. Users can customize these flows by linking to their PDKs, memory compilers, and setting specific constraints. Reports generated through these tools provide power, performance, and area (PPA) estimates. A proof-of-concept SoC using cgra4ml IP integrated into the ARM-based NanoSoC platform demonstrates the viability of this flow.
4.2.4. System Verification
A central component of the backend is the DPI-C-enabled verification suite. Testbenches written in SystemVerilog interface with the C firmware to simulate the complete stack: from quantized model outputs to hardware response. Randomized AXI stimuli stress-test corner cases such as memory congestion or timing edge conditions. This approach reduces the verification burden and ensures fidelity between the model and its physical implementation.
Together, these backend systems allow cgra4ml to bridge the gap between machine learning model development and production-grade hardware deployment, offering flexibility, correctness, and performance portability across platforms.
4.3. Parameterized CGRA Engine Architecture
Figure 7. cgra4ml architecture generated for AMD Xilinx Zynq SoC: Our CGRA-based engine (a) features one AXI-Lite port for configuration and three 128-bit AXI Manager ports for data movement. The CGRA engine (b) performs compute-heavy tasks. The engine has parameterizable bit widths, weights cache, and processing elements (PE) to fit the requirements of different models and meet hardware resource and performance constraints. cgra4ml emits software for the ARM processor (c) to perform pixel-wise operations and handle the edge cases.
cgra4ml architecture generated for AMD Xilinx Zynq SoC
cgra4ml generates a system of hardware components, as described in Fig. 7, that work together with a host processor to process the user defined neural network model. The components are designed to have AXI interfaces for modular hardware design. The system consists of the following subsystems, each with specified parameters that the user sets using the Python API:
  • CGRA - A Coarse Grained Reconfigurable Array (CGRA) with AXI-stream interfaces.
    • Number of Rows & Columns of processing elements (PEs)
    • Bitwidth of inputs, kernel, outputs, and bias
  • Weights cache - an AXI stream module with ping-pong buffers of on-chip SRAM for maximizing data reuse, counters and controllers.
    • Depth of weights cache
    • Maximum batch, kernel sizes, in-channels, height, and width
  • Three Direct Memory Access (DMA) modules that convert between AXI stream and full AXI4.
    • Bitwidth of AXI interfaces (up to 128-bits)
  • Verification IPs and wrappers for full system Verification
    • Valid and Ready probabilities for randomization
In addition to these tunable parameters, clock frequency and I/O delays of the full system are also parametrizable.
cgra4ml implements large and diverse neural networks by creating an efficient yet programmable computational array that can easily be shared across multiple neural network layers. The array’s computation and data movement are design-time parameterizable and run-time programmable to accommodate a variety of DNN layer types. The CGRA is designed to perform various types of neural network layers with optimal performance. This subsection discusses the static, design-time parametrization of the CGRA, while the next subsection discusses runtime configuration.
The static hardware specification defines the CGRA architecture, which can be reprogrammed at runtime within these constraints to execute many different models. These attributes are parameterized into the SystemVerilog hardware to generate different CGRA engines quickly. The parameters are made static before synthesis based on the user-defined attributes. cgra4ml supports fixed-point datatypes, since quantization is critical to meet high throughput and low latency requirements in scientific applications. Their bitwidths are customizable across the inputs, outputs, kernel, and bias. The PE datapaths are automatically adapted to match the requirements of these data types.
The rows and columns dictate the number of PEs in the CGRA, which defines the hardware resource usage and the peak performance. Fig. 7 (d) shows the internal architecture of PEs, which include a multiplier, accumulator, two registers, and two multiplexers. Simplicity is favored over complexity to increase computational density and efficiency. CGRAs designed with a higher degree of flexibility in data movement and operation types suffer from the complexity required to program them and often utilize their resources less effectively. Therefore, our lightweight CGRA design aims to offer minimal programming complexity while supporting high utilization required for scientific edge computing. The PE columns can group themselves and process convolution kernels, sharing data to maximize data reuse.
The depth of the weights cache controls the weight reuse factor. The weights cache consists of two on-chip RAMs with parameterized depths in a ping-pong configuration to allow full throughput. The pixel shifter maximizes data reuse across vertical convolution while maintaining modularity. Their dimensions provide a trade-off between performance and available hardware resources. While any model can be run on any specification, the hardware parameters can be optimized for specific models as described in Subsection 6.1. The maximum limits of layer dimensions determine the register sizes within the hardware (see Table 2).
As shown in Fig. 7, our hardware modules feature open AXI-Full, AXI-Stream, and AXI-Lite interfaces to ensure compatibility with a wide variety of SoCs and other hardware IPs. We utilize open-source AXI and AXI-Stream cores (Alex Forencich, 2024; Alex Forencich., 2024) in our SoC generation, modified to support ASIC flow in addition to FPGA. The AXI interfaces are parameterized to easily match the SoC interfaces. All these modules are wrapped into a single IP that can be easily instantiated and connected to any system with AXI ports.
4.4. Runtime Dynamic Reconfiguration and Dataflow
Table 2. Runtime parameters of the unified dataflow
Description Array Partition
Conv Matmul Slice (S) Iterations (T)
$K_{H}$ Kernel Height 1
$K_{W}$ Kernel Width 1
NN Conv Batch 1
$W$ In Im. Width 1
$H$ In Im. Height I/O Rows $H_{S}{=}R$ $H_{T}{=}\lceil H/H_{S}\rceil$
II In Channels In Cols $I_{S}{=}\lceil D_{W}/K_{H}\rceil$ $I_{T}{=}\lceil I/I_{S}\rceil$
$O$ Out Channels Out Cols $O_{S}{=}\lfloor C/K_{W}\rfloor$ $O_{T}{=}\lceil O/O_{S}\rceil$
  • $D_{W}{=}$ Depth of SRAM in weights rotator
  • R,C=R,C{=} Rows & columns of PE array
Figure 8. Tiling: An output matrix with a shape $[H,O]$ is generated as underlineSlices of shape $H_{S},O_{S}]$ (a), in $HT{\times}OT$ iTerations by a CGRA. The compile-time hardware parameters: depth of weights cache, rows & columns of CGRA ($R{\times}C$) affect the runtime parameters as shown in (b).
Refer to caption Convolution through Unified Dataflow
Figure 9. Convolution through Unified Dataflow: (a) An input tensor of shape [N=1,H=6,W=6,I=3][N{=}1,H{=}6,W{=}6,I{=}3] is convolved with weights tensor of shape $K_{H}{=}5,K_{W}{=}5,I{=}3,O{=}4]$ to generate an output tensor of [N=1,H=6,W=6,O=4][N{=}1,H{=}6,W{=}6,O{=}4]. The runtime parameters (d) are determined by the dimensions of the workload, while the dimensions of the hardware determine the static parameters. In this example configuration, the R,C=4,12R,C{=}4,12 CGRA dimensions (c) map to the slice of the output tensor as shown in (a). The 12 columns of PEs dynamically group into 2 groups of 5 to process the horizontal convolutions ($K_{W}{=}5$). The two groups process two output channels ($O_{S}{=}2$) of the output tensor. The pixel shifter (b) receives 6 words of the input tensor: $R=4$ rows that correspond to the mapping in the output tensor, and $K_{H}/2{=}2$ rows below them. The last $K_{H}/2{=}2$ rows among the $R$ middle rows are stored in an on-chip SRAM to be used as the top rows when processing the next $H_{S}{=}R$ height slice. The top $K_{H}/2{=}2$ rows needed for the current $K_{H}{=5}$ vertical convolution come from this RAM. These $R{+}K_{H}-1{=}8$ words are loaded into a shift register bank and shifted 5 times. The $R$ rows of the CGRA receive the top $R$ rows of the shift register and thus obtain the vertical neighborhood needed for the convolution via shifting. After this shifting, the next input channel of the input image is loaded. After all input channels are loaded and shifted, the PEs of the CGRA (within a group) pass their accumulator values to the PE on their right to be accumulated there. The five PEs in a group (two groups are colored yellow and blue) perform the horizontal convolution this way. Since each group corresponds to an output channel, two such channels are processed in parallel.
In addition to static parametrization of the hardware architecture, cgra4ml also allows dynamic reconfiguration during runtime. Table 2 lists the runtime parameters. cgra4ml’s dataflow and tiling pattern ensures weights are maximally reused, vastly reducing off-chip data movement. The dataflow aims to optimize data reuse patterns; the weights, inputs, and output data move through the CGRA in a predictable pattern for convolutions, dense layers, and matrix multiplications, and can be modified using the cgra4ml attributes. This dataflow is automated, and its control is decentralized across the PEs.
The Python API generates the configuration bits for a layer and passes them to the C firmware. During the setup phase, the firmware writes attributes to the AXI-Lite configuration registers. During the operation, the DMA controller reads the configuration bytes and passes them along with data through the AXI DMAs. The Weights cache and Pixel Shifter receive these configuration bits, process them into fewer bits, and pass them along with the AXI-Streams as TUSER bits to the CGRA. The multiplexers inside the PEs respond to these control bits and reroute the data dynamically at every clock cycle. Since the configuration moves with the data, the CGRA PEs reconfigure themselves on the fly, grouping and sharing partial data to process a layer efficiently, without waiting to clear the pipelines. The runtime parameters and the precise description of our default dataflow through the CGRA are presented in Table 2 and Algorithm 1, respectively.
Next, we describe the two specific cases of dataflow: Matrix Multiplication and Convolution. The flexibility and modularity of the cgra4ml framework enables additional use cases with minimal modification. Additional layers and activation functions are added by modifying the C firmware.
Notation legend. $(\cdot)_{S}$ and $(\cdot)_{T}$ denote slice and iteration partitioning, respectively. We map the $R{\times}C$ PE array to output height and out-channels via $H_{S}{=}R$, $H_{T}{=}\lceil H/H_{S}\rceil$, $O_{S}{=}\lfloor C/K_{W}\rfloor$, $O_{T}{=}\lceil O/O_{S}\rceil$. Input channels are processed with parallel factor $I_{S}$ and iterations $I_{T}{=}\lceil I/I_{S}\rceil$; in our design $I_{S}$ is chosen so $K_{H}\!\cdot\!I_{S}\leq D_{W}$ (weights-SRAM depth), which gives $I_{T}{=}\lceil D_{W}/K_{H}\rceil$ used below. We reorder tensors to expose reuse: $X\!\to\!X_{T}$ (vertical locality via the pixel shifter), $K\!\to\!K_{T}$ (row-major streaming into the weights cache). Braces labeled data beats indicate DMA stream order; those labeled parallel words indicate words consumed/produced per cycle by the array.
$X$ $:[N,H,W,I]$ input
$X_{S}$ $:[N,(H_{T},H_{S}),W,(I_{T},I_{S})]\quad\text{where}\quad H_{T}{=}\lceil H/R\rceil,H_{S}{=}R,I_{T}{=}\lceil D_{W}/K_{H}\rceil,I_{S}{=}\lfloor I/I_{T}\rfloor$ slicing
$X_{T}$ $:\underbrace{[I_{T},N,H_{T},W,I_{S},}_{\text{data beats}}\underbrace{H_{S}{+}K_{H}{-}1]}_{\text{parallel words}}$ reordering
$K$ $:[K_{H},K_{W},I,O]$ weights
$K_{S}$ $:[K_{H},K_{W},(I_{T},I_{S}),(O_{T},O_{S})]\quad\text{where}\quad O_{S}{=}\lfloor C/K_{W}\rfloor,O_{T}{=}\lceil O/O_{S}\rceil$ slicing
$K_{T}$ $:\underbrace{[I_{T},O_{T},}_{\text{(AXI-S Packets)}}\underbrace{I_{S},K_{H},}_{\text{(W-SRAM rows)}}\underbrace{O_{S},K_{W}]}_{\text{(parallel words)}}$ reordering
$Y$ $:[N,H,W,O]$ output
$Y_{S}$ $:[N,(H_{T},H_{S}),W,(O_{T},O_{S})]$ slicing
$Y_{T}$ $:\underbrace{[I_{T},O_{T},N,H_{T},W,}_{\text{data beats}}\underbrace{O_{S},H_{S}]}_{\text{parallel words}}$ reordering
$A^{\prime}$ $:[C,R]$ Accumulators of PE array
$A$ $:[O_{S},K_{W},R]$ dynamic regrouping
$A[:,:,:]\leftarrow 0$
1exfor $i_{t}<I_{T}(=I/I_{S})$ do
 for $o_{t}<O_{T}(=O/O_{S})$ do
    for $n<N$ do
       for $h_{t}<H_{T}(=H/H_{S})$ do
          for $w<W$ do
             for $i_{s}<I_{S}$ do
                 for $k_{h}<K_{H}$ do
                   parallel for $o_{s}<O_{S}(=C/K_{W})$ do
                     parallel for $k_{w}<K_{W}$ do
                       parallel for $h_{s}<H_{S}(=R)$ do
                         $x\leftarrow X_{T}[i_{t},n,h_{t},w,i_{n},h_{s}{+}k_{h}]$
                         $k\leftarrow K_{T}[i_{t},o_{t},i_{s},k_{h},o_{s},k_{w}]$
                         $A[o_{s},k_{w},h_{s}]\mathrel{{+}{=}}k*x$
                         end
                         end
                         end
                         end for
                         end for
[…]
                         parallel for $k_{w}<K_{W}$ do
                           $A[:,k_{w},:]\leftarrow(k_{w}\equiv 0)$ ? $0:A[:,k_{w}{-}1,:]$
                           end
                       end for
                     end for
                   end for
                 end for
Algorithm 1 Unified Dataflow
4.4.1. Matrix Multiplication
Figure 2 shows the tiling (array partitioning), which is primarily output-stationary dataflow. When multiplying two matrices of sizes $[H,I]$ and $[I,O]$ to obtain an output matrix of size $[H,O]$, the output matrix is divided into slices of size $H_{S},O_{S}]$ to be processed in $H_{T}{\times}O_{T}$ iterations, such that $H_{T}{=}\lceil H/H_{S}\rceil$ and $O_{T}{=}\lceil O/O_{S}\rceil$. The rows and columns of the CGRA correspond to the size of a tile that can be processed at a time: $R{\times}C{=}H_{S}{\times}O_{S}$. When processing dense layers, the image batch is presented along $H$ to reuse weights across the batch.
4.4.2. Convolution
The unified dataflow is designed to maximize convolution data reuse. Fig. 9 describes a convolution workload through the unified dataflow with an example workload. $R$ rows of processing elements compute $R{=}H_{S}$ Height slices. Algorithm 1 describes the preparation of the input, weights and output tensors to be processed by the CGRA. The reuse-aware bandwidth model in Eqs. 1-4 of Section 6.1 quantitatively matches the schedule above and guides $R,C,I_{S}$ choices explaining the utilization/data-movement trends in Fig.12.
An input tensor $X$ of shape $[N,H,W,I]$ is sliced according to the static parameters and reshaped into a tensor $X_{T}$ of shape $I_{T},N,H_{T},W,I_{S},R{+}K_{H}{-}1]$. Since the output side has much lower bandwidth, slicing and reshaping the input for the next layer happens at the output side of the current layer, and the words are written to the respective locations in memory to form the tensor $X_{T}$. When processing the next layer, the DMAs read $X_{T}$ in row-major order (as written). The pixel shifter takes $R{+}K_{H}{-}1$ values from the AXI DMA and shifts them, reducing the required input bandwidth by $K_{H}\times$ by exploiting data locality.
The weights cache has two buffers that works in ping-pong configuration, each $C$ words wide and $D_{W}$ rows deep. The weights tensor $K$ of shape $K_{H},K_{W},I,O]$ is sliced and reshaped into a tensor $K_{T}$ of shape $I_{T},O_{T},I_{S},K_{H},O_{S},K_{W}]$. This slicing and reshaping is done offline, by the Python frontend, which then generates the weights binary that is loaded with the firmware, and read by the DMA in row-major order (as is). The weights stored in the weights cache are rotated and reused $O_{T}NH_{T}W$ times, maximizing the weights reuse and reducing the required bandwidth proportionally.
$C$ columns of PEs re-group dynamically into $O_{S}{=}\lfloor C/K_{W}\rfloor$ groups to process one Out-Channel slice $O_{S}$, while each such group of $K_{W}$ columns process the horizontal convolution, exploiting its data locality, and reducing input bandwidth by the same amount. $H_{S}{\times}O_{S}$ output pixels are computed in parallel by the $R{\times}C$ array of processing elements.
Taken together, these components form a robust and extensible foundation for deploying quantized neural networks in embedded scientific systems. By abstracting hardware complexity through modular design and automation, cgra4ml allows scientists and developers to focus on model innovation without being encumbered by low-level engineering overhead. The ability to generate synthesizable RTL, runtime firmware, and verified execution flows ensures consistency across simulation, emulation, and deployment platforms. Moreover, the framework’s flexibility in supporting different processor architectures, SoC shells, and toolchains makes it well-suited to the rapidly evolving demands of scientific edge computing. The seamless integration of verification infrastructure, programmable CGRA engines, and composable bundles underscores the unique value of the framework as a research and deployment tool for ML-powered embedded systems.
5. cgra4ml: Verification and Deployment
Robust verification and deployability are crucial for closing the loop between neural network design and reliable hardware execution, particularly in scientific edge applications, where correctness, reproducibility, and system integration are paramount. While many accelerator frameworks focus on performance or programmability, cgra4ml places equal emphasis on comprehensive verification and streamlined hardware deployment. This section describes the verification methodology that ensures functional consistency across model, firmware, and RTL, followed by the deployment toolflows for FPGA and ASIC targets. Our infrastructure supports agile development and testing through simulation, CI/CD automation, and real-world deployment on programmable SoCs and standard ASIC flows, giving researchers and developers confidence that their designs will perform reliably across the entire development stack.
5.1. Comprehensive Verification Approach
Refer to caption Verification Infrastructure developed for cgra4ml
Figure 10. Verification Infrastructure developed for cgra4ml. The subsystems of CGRA, cache, DMAs and DMA controller are described in SystemVerilog, and during verification are connected to the production-ready C/C++ firmware through AXI-specific VIPs and custom DPI-C wrappers. This way, the same firmware that runs on the FPGA also runs in simulation, verifying the entire system holistically.
Verification is an often overlooked aspect of research designs, which results in significant friction when transitioning from FPGA to ASIC flows. When users implement sophisticated models and developers add new features, it is important to ensure they will work as intended on the final embedded system before investing the effort to implement them. Our verification infrastructure has been tested on Cadence Xcelium, AMD-Xilinx Vivado Xsim, and Verilator (open source).
cgra4ml provides a verification suite that is integrated across hardware, software, and firmware, targeting 100% bit parity with a CPU/GPU golden run at the layer, bundle, and model outputs, using the exact numeric format used by the hardware. Our randomized, transactional testbench suite is built in SystemVerilog and interfaces with the IP’s AXI-Manager ports. The valid and ready probabilities define the randomness in toggling the s_valid and m_ready handshake signals of the AXI interfaces. Setting the probabilities to 1 gives the fastest simulation, where data is transferred every cycle. This is helpful as a smoke test when a user intends to modify the hardware and debug their modification. Setting it to lower values, e.g., 0.1 and 0.01, stress tests the hardware using randomized transactions, exposing hardware bugs that smoke tests cannot reveal. The SystemVerilog Direct Programming Interface for C (DPI-C) is utilized in a novel manner to interface runtime firmware functions with the SV testbench suite to achieve holistic verification.
From the user’s perspective, this verification process is streamlined as follows. First, a randomized tensor is passed through the user’s deep neural network model to capture intermediate and final output tensors in the software frontend. These are sliced, reshaped, and tiled to produce the expected inputs and outputs from the dataflow. During simulation, the data memory resides on the C side and is accessed from the SV side through custom functions. When the simulation begins, a function from the C runtime is called through DPI-C to load the model-specific configuration into the AXI-Lite register bank. As the simulation progresses, the AXI interfaces of the SV testbench read the data memory from the C side and feed the tensors into the RTL design. The outputs from the design are similarly stored into the memory. This process is randomized, such that the valid & ready signals of the address, data and response channels of the AXI interfaces are randomly toggled at the probability set by the user. This randomization emulates the behavior of a congested bus interface, stress-testing the system under realistic conditions and revealing bugs that would otherwise be unidentifiable. The intermediate and final results from the simulation are dumped into files, which the Python API then tests against the expected results.
Our comprehensive verification suite ensures that the model built and trained by the user works as it should on the RTL design and the firmware of the entire embedded system, before even starting the implementation process. All simulations are checked against a CPU/GPU golden reference with identical datatype to the RTL (same data type, per-tensor scale/zero-point, padding/stride/dilation, and the RTL’s rounding and saturation rules). The same checks can be executed on the FPGA build by streaming the identical tensors through the AXI DMA and comparing the captured outputs to the golden results, enabling end-to-end parity checks beyond simulation. After implementation, the same suite can be used on the gate level, time annotated netlist. This allows the developer and the user to add functionality to hardware and the runtime firmware and test it on their desktop before testing it on an FPGA, allowing agile development cycles. A set of GitHub actions running in the cloud verify the entire infrastructure using Verilator for multiple random examples after every update, enabling continuous integration and development (CI/CD), as shown in Figure 6.
5.2. FPGA and ASIC Deployment
cgra4ml is built to generate vendor-agnostic synthesizable SystemVerilog RTL hardware components and necessary scripts to support both FPGA and ASIC implementation flows with various tools and targets.
5.3. FPGA Toolflow
Our design decisions ensure that integrating the generated IP into an FPGA system is fairly straightforward. Modern FPGAs are often Programmable System-on-Chips, where on-chip hardened processors and the FPGA fabric are interconnected to deploy heterogeneous systems. We provide modular TCL scripts to enable such integration, which automates the following process. On AMD-Xilinx FPGAs, the ZYNQ processing system is first configured to have three full AXI Subordinate ports with the maximum width (128-bit on ZCU104) for the highest throughput and one AXI-Lite Manager port for configuration. Their address mappings are updated, and the IP ports are connected to them. The TCL scripts then verify the connections, synthesize, place & route the design, exporting bitstream and generating reports.
In addition to this standard flow, the standardized AXI interfaces of cgra4ml allow the user to easily integrate the generated IP into existing signal processing dataflow systems. Our scripts for AMD-Xilinx Vitis then create an application project, import the bitstream and C firmware, set compiler optimizations, and allow the user to execute the system as baremetal.
PetaLinux, the AMD-specific embedded Linux distribution, is widely used for higher-level applications with ZYNQ systems. When executed on PetaLinux, our C firmware manages the physical to virtual address mapping to provide seamless operation with an intuitive execution API, the same as baremetal. Our C firmware is also wrapped in Python using ctypes, allowing users to pass and receive data as Numpy arrays in a Python environment. This PYNQ API allows the user to be productive during prototyping, similar to the PYNQ API offered by hls4ml. However, as the user moves towards production, our framework also allows layers of the same API to be extended into an optimized production-ready C runtime.
5.4. ASIC Toolflow
ASIC design is a process that often takes several months to years. To obtain a first-order approximation of power-performance-area (PPA) analysis of a given CGRA configuration, we provide a basic set of TCL scripts for Cadence Genus & Synopsys DesignCompiler for synthesis and Cadence Innovus for place & route. Users can link their PDKs, set a few basic parameters, and run the scripts to generate GDSII files and reports. More experienced ASIC designers can further tweak the scripts for a more optimized design.
We have partnered with ARM to tape out an SoC with the IP generated by cgra4ml integrated into NanoSoC (David Flynn, 2023), an open-source SoC platform around an ARM-M0 processor. The platform features APB ports for configuration and three AHB Subordinate ports for data movement in its expansion region. To make our design compatible with their system, we add AXI to AHB bridges around our design and a custom APB Subordinate port in place of the AXI-Lite port.
Together, the verification and deployment flows provided by cgra4ml form a production-grade pipeline that supports model prototyping, debugging, and deployment in a unified environment. By enabling cycle-accurate testing with DPI-C integration, runtime configuration emulation, and full-stack simulation, the framework significantly reduces the time and risk associated with edge ML system development. Furthermore, the modular FPGA and ASIC toolchains allow seamless migration from early experimentation to large-scale integration or even silicon realization. Whether targeting a PYNQ-based research prototype or an industrial-grade SoC, cgra4ml ensures that hardware, software, and machine learning models operate in concert, providing a reliable foundation for next-generation scientific computing at the edge.
5.5. Ibex SoC Integration
To demonstrate end-to-end deployability on an open platform, we integrated cgra4ml with the lowRISC Ibex RISC-V SoC, which we build and simulate via the FuseSoC flow. We add an AXI crossbar to unify the engine’s three AXI managers into a single port, simplifying memory mapping and arbitration. We also add lightweight bridges that translate between the Ibex load/store interface and AXI, allowing the CGRA engine to present a device port for configuration and one host port to the system. The subsystem is then packaged as a reusable FuseSoC IP core. The hardware/firmware interface is platform agnostic. On the software side, we compile our runtime and model firmware with a standard RISC-V toolchain, load the ELF together with quantized weights into on-chip SRAM, and execute the workload on the full SoC. This integration validates cgra4ml ’s hardware/firmware stack in a fully open, reproducible RISC-V environment and provides a template for broader open-source SoC deployments.
6. Experimental Evaluation
In this section, we present the results of experiments comparing cgra4ml to hls4ml on ZCU104 evaluation board with Zynq UltraScale+ XCZU7EV-FFVC1156-2-E MPSoC FPGA. We also demonstrate cgra4ml on deeper models such as ResNet-50 and PointNet, which are too large to be implemented using hls4ml. Our ASIC results show the power and area efficiency of the RTL-based CGRA implemented using the ASIC flow provided with cgra4ml.
6.1. Performance Analysis and Practical Guidelines
The performance of the CGRA for a given bundle can be analyzed based on the runtime parameters ($N,H,H_{S},H_{T},W,I,I_{S},I_{T},O,O_{S},O_{T},K_{W},K_{H}$) described in Table 2. The number of clock cycles and the off-chip data movement required to process a layer on a CGRA with $R,C$ rows & columns of PEs are as follows:
(1) Clock cycles $=O_{T}I_{T}(1+NH_{T}W(1+I_{S}K_{H}))$
(2) Weight words $=O_{T}I_{T}I_{S}K_{H}C$
(3) Input words $=O_{T}I_{T}NH_{T}WI_{S}(R+K_{H}/2)$
(4) Output words $=NH_{\textbf{out}}W_{\textbf{out}}O.$
Increasing the number of PEs improves the performance by increasing the on-chip computation and reducing the number of iterations $O_{T},H_{T}$. The ratio between peak performance ($RC{\times}\text{Frequency}$) and real performance (MAC operations in a layer/time) depends on the ratio of PEs that are idle when computing a layer as follows:
(5) Idle PE cols ratio ${=}[C\%K_{W}]/C{+}[O\%O_{S}]K_{W}/[CO_{T}]$
(6) Idle PE rows ratio ${=}[H{\%}R]/H$
Expression (5) is minimized by our hardware and unified dataflow design (Algorithm 1) that enables the $C$ columns of CGRA to dynamically group and process Out-Channels $O$. For example, for a CGRA with C=96C{=}96, the ratio of idle columns in each iteration: $(C\%K_{W})/C$ is zero for the most common layers with $K_{W}{=}\{1,3\}$, and $1\%,5\%,8\%$ for less common layers with $K_{W}{=}\{5,7,11\}$. To further minimize the overall idle ratio of columns (Exp. 5), the user can pick $C$ such that $\lfloor C/K_{W}\rfloor$ is a factor of $O$ for most layers. For ResNets, multiples of 3 and 4 (C=12,24,96C{=}12,24,96), and for matrix multiplication based workloads, multiples of 8 (C=8,16,32C{=}8,16,32) are such optimal choices. Among the $R$ columns of the CGRA, Expression 6 is minimized by choosing $R$ as the common factor of the $H$ of most layers of the DNNs the user wishes to run on the generated hardware. Figure 12 demonstrates the performance efficiency and off-chip data movement when processing ResNet50 on PE arrays of different sizes.
[…]
Figure 11. Scaling of hardware resource and power consumption on a Xilinx ZCU104 FPGA, for different bitwidths and PE columns, with PE rows = 8
Refer to caption Performance analysis of ResNet-50
Figure 12. Performance analysis of ResNet-50 on different CGRA array sizes. Performance efficiency measures how well each layer is dynamically mapped over statically configured hardware. While larger arrays reduce off-chip data movement due to high data reuse, they have lower utilization for smaller layers. The unified dataflow of the array (Sec. 4.4) maximizes the utilization to near 100% and minimizes data movement for most layers.
Model ResNet-50 PointNet (Qi et al., 2017) Autoencoder (Team, 2025c)
Bits 4 4 8
Operations 5.36 billion 73.7 million 2.09 million
PEs (7,96) (32,32) (16, 64)
Frequency (MHz) 250 250 250
$F$s 101706 69277 154247
LUTs 82200 100076 140357
BRAMs 6 4.5 4.5
Static Power (W) 0.700 0.700 0.702
Dynamic Power (W) 3.847 3.840 4.196
Total Power (W) 4.547 4.540 4.898
GOPs/W 37.3 56.8 1
Table 3. Implementation of ResNet-50 and Pointnet on ZCU104 FPGA
Fig. 11 demonstrates the resource and power consumption for different PE arrays on the FPGA, with rows=8. Since larger arrays are more efficient in terms of GOPs/W, users should strive to implement the largest possible array within the available resources.
Operation Percentage
Writeback pointer calculations 47.37%
Writeback 24.34%
Cache flush write 2.14%
Add IT passes 1.10%
Add Bias 4.50%
Activation 6.04%
Residual add 7.50%
Pooling 4.94%
Handshake 1.18%
Table 4. Breakdown of performance penalty of the CPU workload for ResNet-50
6.2. Performance Evaluation of CPU operations
We partition the workload between the CPU and the CGRA as shown in Fig. 5. Lightweight, pointwise operations are done in the CPU, while heavy, parallelizable operations are done through our CGRA. While this partition allows us implement additional pointwise operations and activations as required, with minimal modification of hardware, our evaluations show that often CPU operations become the bottleneck in addition to data transfer between the CPU and the custom hardware. Depending on the workload, over 50% of the execution time is spent on CPU operations, and Table 4 demonstrates the breakdown of execution time in the CPU for ResNet-50. As observed, we can achieve significant speedup by partially parallelizing the most compute-intensive operations, such as calculating writeback pointers and packing writeback words on custom hardware. We are currently working on this to achieve the performance boost without sacrificing extensibility of the bundle architecture. Our fusion strategy includes writing an extensive set of bundle parameters to a register bank at startup, breaking the CGRA-CPU dual execution into a three-stage flow of CGRA-CPU-Writeback with locks and handshaking for synchronization. The new writeback stage would increase the performance by packing output bits, calculating all pointers in parallel and performing writes without CPU intervention.
6.3. Micro-benchmarking
Refer to caption Synthetic workloads of cascading dense layers
Figure 13. Synthetic workloads of dense autoencoder models with 4-32 layers (x-axis), where each dense layer has 16 inputs and 16 outputs and ReLU activation is used between the layers. The model is quantized to have 8-bit inputs, weights, and biases, which are randomized with a normal distribution. Inference is performed with a batch size of 32. The CGRA is configured to have 16 rows, 32 columns, an accumulator width of 24 bits, and 128-bit wide AXI ports. hls4ml stores weights on-chip due to its dataflow architecture, while cgra4ml stores weights off-chip, but buffers them in ping-pong buffers to reuse across the inference of a layer. We do the benchmarking on the ZCU104 development board. cgra4ml maintains flat resource usage, while hls4ml-generated hardware shows linear growth with the number of layers. We report steady-state inverse-throughput, which is the II (initiation interval) for hls4ml. Latency increases linearly for both, with hls4ml showing a slower rate.
We first conduct a micro-benchmarking experiment to observe how the resource usage and latency of hardware generated by hls4ml and cgra4ml scale with neural network size. The synthetic workload is a stack of dense layers with ReLU between layers; each layer is $16{\times}16$ with bias. All tensors (inputs, weights, biases) are 8-bit quantized; weights and biases are initialized from a zero-mean normal distribution. We report the time to process one batch of 32 inputs. The cgra4ml CGRA uses a $16{\times}32$ PE array with 24-bit accumulators and communicates via 128-bit AXI. On ZCU104, hls4ml stores per-layer weights on-chip (consistent with its layer-specialized dataflow), whereas cgra4ml streams weights from off-chip DDR but employs double (ping–pong) buffering to maximize reuse within each layer. Inputs/outputs are transferred over 128-bit DMA, and layer outputs are requantized to 8-bit with round-to-nearest and saturation. The results are summarized in Fig. 13.
The observations align with our discussion in Section 2. hls4ml, a representative example of layer-by-layer hardware generation, produces designs whose resource usage grows linearly with the number of layers. The hardware dataflow is well pipelined, leading to relatively modest growth in latency. In contrast, cgra4ml is a CGRA-based accelerator that reuses the same core across layers, resulting in flat resource consumption. The latency still grows linearly, but limited pipelining and additional dataflow-control overhead lead to a steeper slope. Based on these trends, we identify three regimes by network depth: for small networks (4–8 layers), cgra4ml achieves lower latency (at higher resource usage) due to its high-throughput systolic array; for mid-sized networks (12–24 layers), hls4ml provides both lower latency and resource usage; beyond 28 layers, cgra4ml becomes attractive because of its constant resource footprint, provided the application can tolerate the higher latency relative to layer-specialized designs.
Figure 14. Iso-Resource Microbenchmarking on Pynq-Z2 at clock frequency of 150 MHz and batch size of 16. For cgra4ml, we find a 16×16 PE array (8×8 MAC, 24-bit accumulator) fills the FPGA. For hls4ml, we find the largest dense layer (8×8 MAC) that can fit within the FPGA with no resources exceeding 100% is that with 32 inputs and 64 outputs (2048 multiply-accumulate operations) at reuse factor of 1. We then increase the workload size linearly by 1024 MACs, and find the reuse factor needed to fit that workload in the FPGA using hls4ml. We find that the reuse factor needs to be increased almost exponentially for linear increase in workload size, causing latency and throughput to degrade for large workload sizes.
To complement Fig. 13, we perform an iso-resource benchmarking on a Pynq-Z2 (150 MHz) with 8-bit activations/weights. For hls4ml, we start with a dense layer of 32 inputs and 64 outputs, which we find is the largest dense layer that fits on the FPGA at reuse factor of 1 without any resources exceeding 100%. The workload size is $32{\times}64{=}2048$ multiply-accumulate operations. We then increase the workload size linearly by 1024, finding the reuse factor necessary to fit it into the FPGA. For CGRA4ML, we select the largest array that fits, which is 16×16 PEs with 8-bit × 8-bit multiply and 24-bit accumulation. In our experiment, we find that LUTs are the bottleneck resource. We run them at the same frequency of 150 MHz and batch size of 16. This way, we keep the resource utilization almost constant (100%) and same for cgra4ml and hls4ml throughout the experiment. We report steady-state batch latency and throughput for both designs in Fig. 14 for every workload size.
Increasing hls4ml reuse is often effective for fitting larger layers. But we find that to fit linearly larger workload size (number of multiply-accumulates), the reuse factor needs to be increased exponentially as shown in Fig. 14. This causes latency and throughput to degrade similarly. At high reuse such as 512, hls4ml latency exceeds that of CGRA4ML, which maintains stable latency and higher throughput at the matched resource point. We also found for input size=128 and output size=64, we could not get resource utilization close to 100% even with reuse factor of 8192, which is the largest possible reuse factor for a dense layer with 8192 MAC operations. Therefore, we suggest the users to prefer hls4ml when the model fits with low reuse factor on the target device, and use CGRA4ML when extreme reuse is required to meet resource limits.
6.4. ResNet-50
ResNet-50 (He et al., 2016) is a deep convolutional neural network with 5.36 billion operations in 53 convolutional layers, one dense layer, two pooling layers, and 17 skip connections. Its input is a $224{\times}224{\times}3$ tensor, and outputs are 1000 classes, and it is run on an array of $7{\times}96{=}672$ PEs.
[…]
Framework cgra4ml hls4ml
Reuse factor - - 1 16 64
PEs (8,12) (16,32) - - -
Frequency (MHz) 250 250 250 250 250
$F$s 32451 64066 21428 20991 21335
LUTs 17959 43320 11810 11986 11963
BRAMs 4.5 4.5 5 5 7
Static Power (W) 0.694 0.697 0.693 0.693 0.693
Dynamic Power (W) 3.078 3.450 2.948 2.886 2.820
Total Power (W) 3.772 4.147 3.642 3.579 3.513
Table 5. Comparing Jet-tagger model on cgra4ml with different array sizes and on hls4ml with varying reuse factors
6.5. PointNet & Jet Tagger
Jet Tagger is a small model with 117K operations in 4 dense layers, widely used for benchmarking. Table 5 compares the Jet Tagger model implemented using cgra4ml with two different CGRA configurations (8×12 and 16×32 PEs) against hls4ml across reuse factors of 1, 16, and 64. Notably, hls4ml achieves lower overall resource usage (FFs and LUTs) and power consumption for this small model. This is expected, as Jet Tagger’s modest compute demand and regular layer structure are well-suited to hls4ml ’s layer-by-layer, statically scheduled architecture, which minimizes hardware overhead when the model fits entirely within the available logic. In contrast, cgra4ml incurs additional overhead from its reconfigurable CGRA infrastructure, runtime firmware, and AXI interfacing, which are optimized for larger, more complex models. For tiny workloads like Jet Tagger, this overhead outweighs the performance benefits of dynamic reconfiguration. Thus, while cgra4ml provides a scalable solution for deeper or irregular networks, hls4ml remains more resource- and power-efficient for simple, compact models, highlighting a natural tradeoff between generality and specialization.
PointNet is a model designed for particle classification. It includes three 1D convolutional layers, three dense layers, and two pooling layers. This results in approximately 1.18 billion operations for a batch size of 16. The model does not synthesize with hls4ml due to its depth, irregular structure, and extensive activation reuse. hls4ml ’s layer-by-layer pipeline architecture requires each layer to be synthesized as a standalone datapath. The entire model must fit into the FPGA fabric simultaneously. Even with reuse factor tuning, hls4ml fails to compile PointNet on the ZCU104 due to resource exhaustion during synthesis and placement. In contrast, cgra4ml supports PointNet by reusing its CGRA fabric across all layers. This enables not only functional deployment but also high energy efficiency (56.8 GOPs/W), making cgra4ml uniquely capable of supporting such mid-sized scientific neural networks that exceed the scalability bounds of traditional HLS-based approaches.
6.6. ASIC Implementation
Technology 7nm 28nm 65nm
Frequency (MHz) 500 500 250
Area (mm2) 0.037 0.175 0.502
Static power (mW) 0.614 25.58 0.130
Dynamic power (mW) 29.8 46.87 80.71
Total power (mW) 30.4 72.45 80.84
GOPs/W 3158 1325 594
Table 6. Power/Performance/Area (PPA) analysis across technology nodes for a CGRA configuration of $R{\times}C=8{\times}24=192$. 28 & 65 nm results use a Cadence Genus/Innovus flow; 7 nm results use a Synopsys DC/ICC2 flow
(a) TSMC 65nm LP (low power)
(b) TSMC 28nm HPC+ (high performance)
(c) TSMC 7nm
(a) 65nm view. (b) 28nm view. (c) 7nm view.
Figure 15. CGRA of $R{\times}C{=}8{\times}24$ configuration generated by cgra4ml and taken through the preliminary ASIC scripts. (a) and (b) are synthesized with Cadence Genus and placed & routed with Cadence Innovus, while (c) was synthesized with Synopsys Design Compiler and placed & routed using Synopsys ICC2.
As described in Sec. 5.4, we built a complete system-on-chip (SoC) around the cgra4ml CGRA using NanoSoC (David Flynn, 2023), an open-source ARM-based SoC generation platform developed with ARM Research. The CGRA connects to the ARM Cortex-M subsystem through DMA360, which bridges the accelerator’s three AXI-Stream ports to the AHB interconnect for programmable, high-throughput transfers. Macros for on-chip memories (weights/pixels) are generated with ARM Artisan Memory Compiler, and the SoC is synthesized in Cadence Genus and placed & routed in Cadence Innovus targeting TSMC 28 nm HPCPLUS. This validates that cgra4ml emits silicon-ready RTL that integrates cleanly into a production-style SoC environment.
Beyond first-order synthesis, we executed place-and-route flows across multiple nodes:
  1. (1)
    TSMC 65 nm LP with Cadence Genus/Innovus
  2. (2)
    TSMC 28 nm HPCPLUS with Cadence Genus/Innovus, and
  3. (3)
    TSMC 7 nm with Synopsys DC/ICC2
For 65 nm, we targeted 250 MHz with a custom floorplan that partitions the weight memory into narrow (8-bit) SRAM macros placed symmetrically near AXI-Stream port for weights, and a 12-bit macro for pixels; the resulting design uses 9 SRAM macros. At 28 nm and 7 nm, the SRAM instances are sufficiently compact to avoid width-splitting; these designs employ 3 macros total. Across nodes we used scripted flows from synthesis through P&R, including memory macro integration, floorplanning, clocking, routing, and iterative DRC cleanup.
In addition to the 65 nm and 28 nm implementations above, we evaluated larger cgra4ml-generated arrays in TSMC 7 nm. The 768-PE configuration closes at 1 GHz with sub-0.5 W total power (Table 6), illustrating the scalability of cgra4ml toward high-throughput ASIC targets. By varying $R{\times}C$ and memory organization from the Python frontend, users can rapidly explore power, performance, and area trade-offs across process nodes and CGRA sizes.
These results validate that the RTL emitted by cgra4ml is not tied to any single foundry or node, but can be retargeted across a wide spectrum of technologies. The ability to support both bleeding-edge 7nm and legacy 65nm nodes makes cgra4ml suitable for deployment in academic tapeouts, low-cost SoCs, and commercial research settings alike. By varying the CGRA configuration parameters from the Python frontend, users can easily explore the power, performance, and area trade-offs across process technologies. This highlights the practical utility of cgra4ml not only as a research framework but also as a prototyping and pre-silicon exploration tool for real ASIC targets. The scripts and toolflows we provide are for preliminary exploration. Due to their nature each tapeout requires an experienced physical designer to extensively tweak the scripts and flow to generate results free of DRC violations. cgra4ml allows rapid iteration on ASIC-ready CGRA accelerators, from low-power edge systems to high-performance scientific computing chips.
7. Future Work
We are expanding the framework’s capabilities to support more complex neural network architectures, including transformers, to broaden its applicability in scientific edge computing further. We are also working on migrating more and more runtime components to hardware, to parallelize their computation and minimize any CPU bottlenecks.
8. Conclusion
In this paper, we introduced cgra4ml, a modular framework for deploying large-scale neural networks in scientific edge computing. By supporting off-chip storage, diverse model architectures, and holistic ASIC/FPGA flows, cgra4ml complements existing tools like hls4ml while enabling more complex and performant designs. Its vendor-agnostic RTL generation, runtime firmware, and integrated verification infrastructure reduce development overhead and simplify hardware-software co-design. We believe cgra4ml offers a practical and extensible foundation for advancing high-performance scientific computing at the edge.